### v5

# A simpler and parallelizable $O(\sqrt{\log n})$-approximation algorithm for Sparsest Cut

Abstract
Currently, the best known tradeoff between approximation ratio and complexity for the Sparsest Cut problem is achieved by the algorithm in [Sherman, FOCS 2009]: it computes $O(\sqrt{(\log n)/\varepsilon})$-approximation using $O(n^{\varepsilon}\log^{O(1)}n)$ maxflows for any $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$. It works by solving the SDP relaxation of [Arora-Rao-Vazirani, STOC 2004] using the Multiplicative Weights Update algorithm (MW) of [Arora-Kale, JACM 2016]. To implement one MW step, Sherman approximately solves a multicommodity flow problem using another application of MW. Nested MW steps are solved via a certain “chaining” algorithm that combines results of multiple calls to the maxflow algorithm.
We present an alternative approach that avoids solving the multicommodity flow problem and instead computes “violating paths”. This simplifies Sherman’s algorithm by removing a need for a nested application of MW, and also allows parallelization: we show how to compute $O(\sqrt{(\log n)/\varepsilon})$-approximation via $O(\log^{O(1)}n)$ maxflows using $O(n^{\varepsilon})$ processors.
We also revisit Sherman’s chaining algorithm, and present a simpler version together with a new analysis.
1 Introduction
Partitioning a given undirected graph $G=(V,E)$ into two (or more) components is a fundamental problem in computer science with many real-world applications, ranging from data clustering and network analysis to parallel computing and VLSI design. Usually, desired partitions should satisfy two properties: (i) the total cost of edges between different components should be small, and (ii) the components should be sufficiently balanced. For partitions with two components $(S,\bar{S})$ this means that $E(S,\bar{S})$ should be small and $\min\{|S|,|\bar{S}|\}$ should be large, where $E(S,\bar{S})$ is the total number of edges between $S$ and $\bar{S}$ (or their total weight in the case of weighted graphs).
One of the most widely studied versions is the Sparsest Cut problem whose goal is to minimize the ratio $\frac{E(S,\bar{S})}{\min\{|S|,|\bar{S}|\}}$ (called edge expansion) over partitions $(S,\bar{S})$. Another well-known variant is the $c$-Balanced Separator problem: minimize $E(S,\bar{S})$ over $c$-balanced partitions, i.e. partitions satisfying ${\min\{|S|,|\bar{S}|\}}\geq cn$ where $n=|V|$ and $c\in(0,\tfrac{1}{2})$ is a given constant.
Both problems are NP-hard, which forces one to study approximation algorithms, or pseudoapproximation algorithms in the case of Balanced Separator. (An algorithm for Balanced Separator is said to be a $\lambda$-pseudoapproximation if it computes a $c^{\prime}$-balanced partition $(S,\bar{S})$ whose cost $E(S,\bar{S})$ is at most $\lambda$ times the optimal cost of the $c$-Balanced Partition problem, for some constants $c^{\prime}\leq c$). Below we discuss known results for the Sparsest Cut problem. They also apply to Balanced Separator: in all previous works, whenever there is an $O(\lambda)$-approximation algorithm for Sparsest Cut then there is an $O(\lambda)$-pseudoapproximation algorithm for Balanced Separator with the same complexity.
The first nontrivial guarantee was obtained by Leighton and Rao [LR99], who presented a $O(\log n)$-approximation algorithm based on a certain LP relaxation of the problem. The approximation factor was improved to $O(\sqrt{\log n})$ in another seminal paper by Arora, Rao and Vazirani [ARV09] who used an SDP relaxation. Arora, Hazan and Kale [AHK10] showed how to (approximately) solve this SDP in $\tilde{O}(n^{2})$ time using multicommodity flows while preserving the $O(\sqrt{\log n})$ approximation factor. Arora and Kale later developed in [AK16] a more general method for solving SDPs that allowed different tradeoffs between approximation factor and complexity; in particular, they presented an $O(\log n)$-approximation algorithm using $O(\log^{O(1)}n)$ maxflow computations, and a simpler version of $O(\sqrt{\log n})$-approximation with $\tilde{O}(n^{2})$ complexity.
The algorithms in [AHK10, AK16] were based on the Multiplicative Weights Update method. An alternative approach based on the so-called cut-matching game was proposed by Khandekar, Rao and Vazirani [KRV06]; their method computes $O(\log^{2}n)$-approximation for Sparsest Cut using $O(\log^{O(1)}n)$ maxflows. This was later improved to $O(\log n)$-approximation by Orecchia, Schulman, Vazirani and Vishnoi [OSVV08].
The line of works above culminated in the result of Sherman [She09], who showed how to compute $O(\sqrt{(\log n)/\varepsilon})$-approximation for Sparsest Cut using $O(n^{\varepsilon}\log^{O(1)}n)$ maxflows for any $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$. This effectively subsumes previous results, as taking $\varepsilon=\Theta(1/\log n)$ yields an $O(\log n)$ approximation using $O(\log^{O(1)}n)$ maxflows, while a sufficiently small constant $\varepsilon$ achieves an $O(\sqrt{\log n})$-approximation and improves on the $\tilde{O}(n^{2})$ runtime in [AHK10, AK16]. In particular, using the recent almost linear-time maxflow algorithm [CKL+22] yields $O(n^{1+\varepsilon})$ complexity for $O(\sqrt{\log n})$-approximation. (As usual, we assume in this paper that graph $G$ has $m=O(n\log n)$ edges. This can be achieved by sparsifying the graph using the algorithm of Benczúr and Karger [BK96], which with high probability preserves the cost of all cuts up to any given constant factor.)
Our contributions   In this paper we present a new algorithm that computes an $O(\sqrt{(\log n)/\varepsilon})$ approximation for Sparsest Cut w.h.p. whose expected runtime is $O((n^{\varepsilon}\log^{O(1)}n)\cdot T_{\tt maxflow})$ for given $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$, where $T_{\tt maxflow}=\Omega(n)$ is the runtime of a maxflow algorithm on a graph with $n$ nodes and $O(n\log n)$ edges. It has the following features:
(i) It simplifies Sherman’s algorithm in two different ways.
(ii) The algorithm is parallelizable: it can be implemented on $O(n^{\varepsilon})$ processors in expected parallel runtime $O((\log^{O(1)}n)\cdot T_{\tt maxflow})$ (in any version of the PRAM model), where “parallel runtime” is defined as the maximum runtime over all processors.
(iii) To prove algorithm’s correctness, we introduce a new technique, which we believe may yield smaller constants in the $O(\cdot)$ notation. Note that there are numerous papers that optimize constants for problems with a constant factor approximation guarantee. We argue that this direction makes just as much sense for the Sparsest Cut problem. The question can be naturally formulated as follows: what is the fastest algorithm to compute $C\sqrt{\log n}$-approximation for a given constant $C$? Alternatively, for maxflow-based algorithms one may ask what is the smallest $C=C_{\varepsilon}$ such that there is a $C\sqrt{\log n}$-approximation algorithm that uses $\tilde{O}(n^{\varepsilon})$ maxflow, for given $\varepsilon>0$. Unfortunately, it is impossible to directly compare our constant with that of Sherman: we believe that the paper [She09] contains a numerical mistake (see the footnote in Section 3.4). An additional complication is that optimizing the constants may not be an easy task. Due to these considerations, we formulate our claim differently: our proof technique should lead to smaller constants since our analysis is more compact and avoids case analysis present in [She09].
To explain details, we need to give some background on Sherman’s algorithm. It builds on the work of Arora and Kale [AK16] who approximately solve an SDP relaxation using a (matrix) Multiplicative Weights update algorithm (MW). The key subroutine is to identify constraints that are violated by the current primal solution. Both [AK16] and [She09] do this by solving a multicommodity flow problem. Arora and Kale simply call Fleischer’s multicommodity flow algorithm [Fle00], while Sherman designs a more efficient customized method for approximately solving this flow problem using another application of MW. Our algorithm avoids solving the multicommodity flow problem, and instead searches for “violating paths”, i.e. paths that violate triangle inequalities in the SDP relaxation. We show that this can be done by a simple randomized procedure that does not rely on MW. Furthermore, independent calls to this procedure return violating paths that are mostly disjoint, which allows parallelization: we can compute many such paths on different processors and then take their union.
In order to compute violating paths, we first design procedure ${\tt Matching}(u)$ that takes vector $u\in\mathbb{R}^{d}$ and outputs a directed matching on a given set $S\subset\mathbb{R}^{d}$ with $|S|=\Theta(n)$. (It works by calling a maxflow algorithm and then postprocessing the flow). The problem then boils down to the following: given vector $u$ randomly sampled from the Gaussian distribution, we need to sample vectors $u_{1},\ldots,u_{K}$ so that set ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{K})$ contains many paths $(x_{0},x_{1},\ldots,x_{K})$ with “large stretch”, i.e. with $\langle x_{K}-x_{0},u\rangle\geq\Omega(K)$. (Here $\circ$” is the operation that “chains together” paths in a natural way). This was also a key task in [She09], where it was needed for implementing one inner MW step. We refer to an algorithm that samples vectors $u_{1},\ldots,u_{K}$ as a chaining algorithm.
Sherman’s chaining algorithm can be viewed as an algorithmization of the proof in the original ARV paper [ARV09] and its subsequent improvement by Lee [Lee05]. We present a simpler chaining algorithm with a very different proof; as stated before, the new proof may yield smaller constants.
Concurrent work   After finishing the first draft of the paper [Kol23], we learned about a very recent work by Lau-Tung-Wang [LTW24], which considers a generalization of Sparsest Cut to directed graphs. The authors presented an algorithm which, in the case of undirected graphs, also simplifies Sherman’s algorithm by computing violating paths instead of solving a multicommodity flow problem. Unlike our paper, [LTW24] does not consider parallelization, and uses Sherman’s chaining algorithm as a black box.
Another very recent paper by Agarwal-Khanna-Li-Patil-Wang-White-Zhong [AKL+24] presented a parallel algorithm for approximate maxflow with polylogarithmic depth and near-linear work in the PRAM model. Using this algorithm, they presented, in particular, an $O(\log^{3}n)$-approximation sparsest cut algorithm with polylogarithmic depth and near-linear work (by building on the work [NS17] who showed how to replace exact maxflow computations in the cut-matching game [KRV06, OSVV08] with approximate maxflows).
Other related work   The problem of computing a cut of conductance $\tilde{O}(\sqrt{\phi})$ assuming the existence of a cut of conductance $\phi\in[0,1]$ (with various conditions on the balancedness) has been considered in [SMP15, KM15, OV11]. Papers [SMP15, KM15] presented distributed algorithms for this problem (in the CONGEST model), while [PS14] observed that the BalCut algorithm in [OV11] is parallelizable: it can be implemented in near-linear work and polylogarithmic depth. Note that $\phi$ can be much smaller than 1. [CS19] presented a distributed CONGEST algorithm for computing an expander decomposition of a graph.
2 Background: Arora-Kale framework
We will describe the algorithm only for the $c$-Balanced Separator problem. As shown in [AK16], the Sparsest Cut problem can be solved by a very similar approach, essentially by reducing it to the $c$-Balanced Separator problem for some constant $c$; we refer to [AK16] for details.111 These details are actually not included in the journal version [AK16], but can be found in [Kal07] or in https://www.cs.princeton.edu/~arora/pubs/mmw.pdf, Appendix A.
Let $c_{e}\geq 0$ be the weight of edge $e$ in $G$. The standard SDP relaxation of the $c$-Balanced Separator problem can be written in vector form as follows (see [ARV09]):
$\min\sum_{e=\{i,j\}\in E}$ $c_{e}||v_{i}-v_{j}||^{2}$ (1a)
$||v_{i}||^{2}$ $=1$ $\forall i$ (1b)
$||v_{i}-v_{j}||^{2}+||v_{j}-v_{k}||^{2}$ $\geq||v_{i}-v_{k}||^{2}$ $\forall i,j,k$ (1c)
$\sum_{i<j}||v_{i}-v_{j}||^{2}$ $\geq 4c(1-c)n^{2}$ (1d)
Here $v_{i}\in\mathbb{R}^{n}$ for each node $i\in V$. The optimum of this SDP divided by 4 is a lower bound on the minimum $c$-Balanced Separator problem. Arora and Kale considered a slightly different relaxation:
$\min\sum_{e=\{i,j\}\in E}$ $c_{e}||v_{i}-v_{j}||^{2}$ $\min C$ $\bullet X$ (2a)
$||v_{i}||^{2}$ $=1$ $X_{ii}$ $=1$ $\forall i$ (2b)
$\sum_{j=1}^{\ell(p)}||v_{p_{j}}-v_{p_{j-1}}||^{2}$ $\geq||v_{p_{\ell(p)}}-v_{p_{0}}||^{2}$ $T_{p}\bullet X$ $\geq 0$ $\forall p$ (2c)
$\sum_{i,j\in S:i<j}||v_{i}-v_{j}||^{2}$ $\geq\xi n^{2}$ $K_{S}\bullet X$ $\geq\xi n^{2}$ $\forall S$ (2d)
$X$ $\succeq 0$ (2e)
Here $p$ stands for a path $p=(p_{0},\ldots,p_{\ell(p)})$ in graph $G$, notation “$\forall S$” means all subsets $S\subseteq V$ of size at least $(1-c/4)n$, and $\xi=3c-4c^{2}$. Matrix $X$ is defined via $X=V^{T}V$ where $V$ is the $n\times n$ matrix with columns $v_{1},\ldots,v_{n}$. We have $C={\cal L}(G)$ and $T_{p}={\cal L}(G_{p})-{\cal L}(G_{(p_{0},p_{\ell(p)})})$ where ${\cal L}(\cdot)$ is the Laplacian of the corresponding graph and $G_{q}$ for a path $q$ is the undirected unweighted graph containing all edges of $q$. Finally, $K_{S}(i,i)=|S|-1$ for $i\in S$, $K_{S}(i,j)=-1$ for distinct $i,j\in S$, and $K_{S}(i,j)=0$ in all other cases.
Note that triangle inequalities (1c) imply path inequalities (2c), while constraints (1b) and (1d) imply constraints (2d) (see [AK16]). SDP (2) may be looser than (1), but its optimum divided by 4 is still a lower bound on the minimum $c$-Balanced Separator problem.
In the sequel we will use a slight modification of (2) in which constraints (2c) are enforced for all sequences $p=(p_{0},\ldots,p_{\ell(p)})$ of distinct nodes in $G$, and not just paths in $G$. This modification can only make the relaxation stronger; constraints (2c) are now equivalent to the triangle inequalities in (1c).
The dual of (2) is as follows. It has variables $y_{i}$ for every node $i$, $f_{p}$ for every path $p$, and $z_{S}$ for every set $S$ of size at least $(1-c/4)n$. Let ${\tt diag}(y)$ be the diagonal matrix with vector $y$ on the diagonal:
$\max\sum_{i}y_{i}+\xi n^{2}\sum_{S}z_{S}$ (3a)
${\tt diag}(y)+\sum_{p}f_{p}T_{p}+\sum_{S}z_{S}K_{S}\preceq C$ (3b)
$f_{p},z_{S}\geq 0\quad\forall p,S$ (3c)
2.1 Matrix multiplicative weights algorithm
To solve the above SDP, [AK16] converts it to a sequence of feasibility problems. First, an interval $[L,U]$ is computed containing an optimal value of the objective. One can use, for example, the algorithm in [OSVV08] to get such an interval with $U/L=O(\log n)$ via $O(\log^{O(1)}n)$ maxflows. Fix a value $\alpha\in[L,U]$, and replace the objective (2a) with the constraint
$C\bullet X\leq\alpha$ (2a)
It suffices to try $O(\log(U/L))=O(\log\log n)$ values of threshold $\alpha$ if we are willing to accept the loss by a constant factor in the approximation ratio.
[…]
Input: vectors $v_{1},\ldots,v_{n}\in\mathbb{R}^{n}$ and $\tilde{v}_{1},\ldots,\tilde{v}_{n}\in\mathbb{R}^{d}$ satisfying $||\tilde{v}_{i}||^{2}$   $\;\;\leq\;\;2$ $\forall i$ (4a) $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$ $\;\;\geq\;\;\tfrac{\xi n^{2}}{4}$ (4b) $|\;||\tilde{v}_{i}||^{2}-||v_{i}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}||^{2}+\tau)$ $\forall i$ (4c) $|\;||\tilde{v}_{i}-\tilde{v}_{j}||^{2}-||v_{i}-v_{j}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}-\tilde{v}_{j}||^{2}+\tau)$ $\forall i,j$ (4d) for some constants $\gamma,\tau>0$. Let $X=V^{T}V$ and $\tilde{X}=\tilde{V}^{T}\tilde{V}$ where $V\in\mathbb{R}^{n\times n}$ and $\tilde{V}\in\mathbb{R}^{d\times n}$ are the matrices with columns $\{v_{i}\}$ and $\{\tilde{v}_{i}\}$, respectively. Output: either (i) variables $f_{p}\geq 0$ and symmetric matrix $F\preceq C$ such that $\left(\sum\nolimits_{p}f_{p}T_{p}-F\right)\bullet X\leq-\alpha$ (5) or (ii) a $\Theta(1)$-balanced cut of value at most $\kappa\alpha$.
The number of iterations of the MW algorithm will depend on the maximum possible spectral norm of matrix $\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$. This parameter is called the width of the oracle, and will be denoted as $\rho$. We will use the bound $\rho=||\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F||\leq\tfrac{\alpha}{n}+||\sum_{p}f_{p}T_{p}-F||$. The algorithm has the following guarantee (see Appendix A for details).
Theorem 1 ([AK16]).
If in the MW algorithm the first $T=\lceil\frac{4\rho^{2}n^{2}\ln n}{\epsilon^{2}}\rceil$ calls to Oracle output option (i) then the optimum value of SDP (2) is at least $\alpha-\epsilon$.
2.2 Oracle implementation
To implement the oracle, Arora and Kale interpret values $f_{p}$ as a multicommodity flow in graph $G$, i.e. a flow that sends $f_{p}$ units of demand between the endpoints of $p$. Given this flow, introduce the following notation. Let $f_{e}$ be the flow on edge $e$ (i.e. $f_{e}=\sum_{p\ni e}f_{p}$). Let $d_{ij}$ be the total flow between nodes $i$ and $j$ (i.e. $d_{ij}=\sum_{p\in{\cal P}_{ij}}f_{p}$ where ${\cal P}_{ij}$ is the set of paths from $i$ to $j$). Finally, let $d_{i}$ be the total flow from node $i$ (i.e. $d_{i}=\sum_{j}d_{ij})$. Given parameter $\pi>0$, a valid $\pi$-regular flow is one that satisfies capacity constraints: $f_{e}\leq c_{e}$ for all edges $e$ and $d_{i}\leq\pi$ for all nodes $i$.
The oracle in [AK16] computes a $\pi$-regular flow $f$ for some parameter $\pi$, and sets $F$ to be the Laplacian of the flow graph (i.e. the weighted graph where edge $e$ has weight $f_{e}$). Capacity constraints then ensure that $F\preceq C$ (because $C-F$ is the Laplacian of the weighted graph with weights $c_{e}-f_{e}\geq 0$ on edges $e\in E$). Let $D$ be the Laplacian of the demand graph (i.e. the complete weighted graph where edge $\{i,j\}$ has weight $d_{ij}$). It can be checked that $\sum_{p}f_{p}T_{p}=F-D$. Thus, the oracle needs to ensure that $D\bullet X\geq\alpha$, or equivalently
$\sum_{i<j}d_{ij}||v_{i}-v_{j}||^{2}\geq\alpha$ (6)
All degrees in the demand graph are bounded by $\pi$, therefore $||D||\leq 2\pi$. Thus, the width of the oracle can be bounded as $||\rho||\leq\tfrac{\alpha}{n}+||D||\leq\tfrac{\alpha}{n}+2\pi$.
Below we summarize three known implementations of the oracle. The first two are due to Arora and Kale [AK16] and the third one is due to Sherman [She09].
  1. 1.
    Using $O(1)$ expected maxflow computations, the oracle computes either a $\pi$-regular flow with $\pi=O(\tfrac{\alpha\log n}{n})$ or a $\Theta(1)$-balanced cut of capacity at most $O(\alpha\log n)$.
  2. 2.
    Using $O(1)$ expected multicommodity flow computations, the oracle computes either a $\pi$-regular flow with $\pi=O(\tfrac{\alpha}{n})$ or a $\Theta(1)$-balanced cut of capacity at most $O(\alpha\sqrt{\log n})$.
  3. 3.
    Let $\varepsilon\in[O(1/\log n),\Omega(1)]$. Using $O(n^{\varepsilon}\log^{O(1)}n)$ expected maxflow computations, the oracle either computes a $\pi$-regular flow with $\pi=O(\tfrac{\alpha}{\varepsilon n})$, or a $\Theta(1)$-balanced cut of capacity at most $O\left(\alpha\sqrt{\tfrac{\log n}{\varepsilon}}\right)$.
By the discussion in Section 2.1, these oracles lead to algorithms with approximation factors $O(\log n)$, $O(\sqrt{\log n})$ and $O\left(\sqrt{\tfrac{\log n}{\varepsilon}}\right)$, respectively.
To conclude this section, we discuss how to verify condition (6) in practice. (Recall that we only have an access to approximations $\tilde{v}_{i}$ of vectors $v_{i}$.)
Proposition 2.
Suppose parameters $\tau,\gamma$ in eq. (4c)-(4d) satisfy $\tau\leq 2$ and $\gamma\leq\tfrac{\alpha}{20n\pi}$. Then condition
$\sum_{i<j}d_{ij}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}\geq 2\alpha$ (7)
implies condition (6).
Proof.
Denote $z_{ij}=||v_{i}-v_{j}||^{2}$ and $\tilde{z}_{ij}=||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$. We have $||\tilde{v}_{i}||^{2}\leq 2$, $||\tilde{v}_{j}||^{2}\leq 2$ and hence $\tilde{z}_{ij}\leq 8$. By Theorem 21 we then have $|\tilde{z}_{ij}-z_{ij}|\leq\gamma(\tilde{z}_{ij}+\tau)\leq 10\gamma$. This implies that $\sum_{i<j}d_{ij}|\tilde{z}_{ij}-z_{ij}|\leq 10\gamma\cdot\sum_{i<j}d_{ij}\leq 10\gamma\cdot 2n\pi\leq\alpha$. The claim follows. ∎
3 Our algorithm
In this section we present our implementation of the oracle. To simplify notation, we assume in this section that vectors $\tilde{v}_{i}$ for $i\in V$ are unique, and rename the nodes in $V$ so that $\tilde{v}_{x}=x$ for each $x\in V$. Thus, we now have $V\subseteq\mathbb{R}^{d}$. The “true” vector in $\mathbb{R}^{n}$ corresponding to $x\in V$ is still denoted as $v_{x}$.
Recall that the oracles in [AK16, She09] do one of the following:
  • output a cut;
  • output multicommodity flows $f_{p}$ satisfying (6), and set $F$ to be its flow graph.
Our oracle will use a third option described in the lemma below.
Lemma 3.
Let $M$ be a set of paths on $V$ such that each $p\in M$ violates the path inequality by some amount $\tfrac{1}{2}\Delta>0$. In other words, we require that $T_{p}\bullet X\leq-\tfrac{1}{2}\Delta$, or equivalently
$\sum_{j=1}^{\ell(p)}||v_{p_{j}}-v_{p_{j-1}}||^{2}\;\;\leq\;\;||v_{p_{\ell(p)}}-v_{p_{0}}||^{2}-\tfrac{1}{2}\Delta$ (8)
Let ${\cal G}_{F}$ and ${\cal G}_{D}$ be respectively flow and demand graphs of the multicommodity flow defined by $M$ (where each path carries one unit of flow). Set $f_{p}=\frac{2\alpha}{|M|\Delta}$ for all $p\in M$, $f_{p}=0$ for $p\notin M$, and $F=0$. Then these variables give a valid output of the oracle with width $\rho\leq\tfrac{\alpha}{n}+\tfrac{4\alpha(\pi_{F}+\pi_{D})}{|M|\Delta}$ where $\pi_{F},\pi_{D}$ are the maximum degrees of ${\cal G}_{F},{\cal G}_{D}$, respectively.
Proof.
We have $F\preceq C$ and $(\sum_{p}T_{p}-F)\bullet X\leq\sum_{p}f_{p}\cdot(-\tfrac{1}{2}\Delta)=-\alpha$, so condition (5) holds. It can be checked that $\sum_{p}f_{p}T_{p}=\frac{2\alpha}{|M|\Delta}(\tilde{F}-\tilde{D})$ where $\tilde{F}$ and $\tilde{D}$ are the Laplacians of respectively ${\cal G}_{F}$ and ${\cal G}_{D}$. Therefore, $\rho=||\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F||=||\tfrac{\alpha}{n}I+\tfrac{2\alpha}{|M|\Delta}(\tilde{F}-\tilde{D})||\ \leq||\tfrac{\alpha}{n}I||+\tfrac{2\alpha}{|M|\Delta}(||\tilde{F}||+||\tilde{D}||)\leq\tfrac{\alpha}{n}+\tfrac{2\alpha}{|M|\Delta}(2\pi_{F}+2\pi_{D})$. ∎
Recall that in the specification of the oracle we required paths $p$ to have distinct nodes (to make the number of constraints finite). This does not cause problems for Lemma 3: if some path in $M$ does not satisfy this, then before applying the lemma we can shorten it while preserving endpoints and condition (8).
The following proposition shows how to verify condition (8) for unobserved variables $v_{x}\in\mathbb{R}^{n}$ using observed variables $x\in\mathbb{R}^{d}$. Its proof is very similar to that of Proposition 2, and is omitted.
Proposition 4.
Suppose parameters $\tau,\gamma$ in Theorem 21 satisfy $\tau\leq 2$ and $\gamma\leq\tfrac{\Delta}{20(K+1)}$ for some integer $K\geq 1$. Then condition
$\sum_{j=1}^{\ell(p)}||p_{j}-p_{j-1}||^{2}\;\;\leq\;\;||p_{\ell(p)}-p_{0}||^{2}-\Delta$ (9)
implies condition (8), assuming that $\ell(p)\leq K$.
Remark 1.
Sherman [She09] explicitly tries to (approximately) solve the multicommodity flow problem: find valid flows $\{f_{p}\}$ in $G$ with demands $\{d_{xy}\}$ that maximize $\sum_{x,y}d_{xy}||x-y||^{2}$. This is done by an iterative scheme via the Multiplicative Weights (MW) framework. Using the option in Lemma 3 has the following advantages over this approach.
  1. 1.
    We can avoid another application of MW and thus simplify the algorithm.
  2. 2.
    The oracle can be easily parallelized: we can compute different “violating paths” on different processors and then take their union. (Of course, we still need to make sure that these paths are “sufficiently disjoint” so that the degrees of graphs ${\cal G}_{F},{\cal G}_{D}$ and hence the oracle width remain small).
3.1 Correlated Gaussians and measure concentration
We write $u\sim{\cal N}$ to indicate that $u$ is a random vector in $\mathbb{R}^{d}$ with Gaussian independent components $u_{i}\sim{\cal N}(0,1)$. Throughout the paper notation ${\tt Pr}_{u}[\cdot]$ means the probability under distribution $u\sim{\cal N}$. We write $(u,u^{\prime})\sim{\cal N}_{\omega}$ for $\omega\in[0,1)$ to indicate that $(u,u^{\prime})$ are random vectors in $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that for each $i\in[d]$, pair $(u_{i},u^{\prime}_{i})$ is an independent 2-dimensional Gaussian with mean $(0,0)^{T}$ and covariance matrix $\begin{pmatrix}1&\omega\\ \omega&1\end{pmatrix}$. We write $u^{\prime}\sim_{\omega}u$ to indicate that $u^{\prime}$ is an $\omega$-correlated copy of $u$ [She09], i.e. $(u,u^{\prime})$ is generated according to $(u,u^{\prime})\sim{\cal N}_{\omega}$ conditioned on fixed $u$. It can be checked that for each $i\in[d]$, $u^{\prime}_{i}$ is an independent Gaussian with mean $\omega\cdot u_{i}$ and variance $1-\omega^{2}$. Note, if $(u,u^{\prime})\sim{\cal N}_{\omega}$ then $u\sim{\cal N}$ and $u^{\prime}\sim{\cal N}$. Conversely, the process $u\sim{\cal N},u^{\prime}\sim_{\omega}u$ generates pair $(u,u^{\prime})$ with distribution ${\cal N}_{\omega}$. The same is true for the process $u^{\prime}\sim{\cal N},u\sim_{\omega}u^{\prime}$.
The key property for obtaining an $O(\sqrt{\log n})$-approximation algorithm is measure concentration of the Gaussian distribution. This property can be expressed in a number of different ways; we will use the following version. [^2] Theorem 5 is formulated in [MOR+06] for the discrete cube, but the proof also works for the Gaussian distribution. For completeness, we reproduce the proof in Appendix B.
Theorem 5 ([MOR+06]).
Consider sets ${\cal A}\subseteq\mathbb{R}^{d}$, ${\cal B}\subseteq\mathbb{R}^{d}$ with ${\tt Pr}_{u}[u\in{\cal A}]={\tt Pr}_{u^{\prime}}[u^{\prime}\in{\cal B}]=\delta$. Then
${\tt Pr}_{(u,u^{\prime})\sim{\cal N}_{\omega}}[(u,u^{\prime})\in{\cal A}\times{\cal B}]\geq\delta^{2/(1-\omega)}$
Given a sequence of numbers $\omega_{1},\ldots,\omega_{k-1}\in[0,1)$, we write $(u_{1},\ldots,u_{k})\sim{\cal N}_{\omega_{1},\ldots,\omega_{k-1}}$ to indicate the following distribution: sample $u_{1}\sim{\cal N}$, then $u_{2}\sim_{\omega_{1}}u_{1}$, then $u_{3}\sim_{\omega_{2}}u_{2}$, $\ldots$, then $u_{k}\sim_{\omega_{k-1}}u_{k-1}$. If $\omega_{1}\!=\!\ldots\!=\!\omega_{k-1}\!=\!\omega$ then we write ${\cal N}^{k}_{\omega}$ instead of ${\cal N}_{\omega_{1},\ldots,\omega_{k-1}}$ for brevity. Finally, if some values of the sequence $(u_{1},\ldots,u_{k})$ are fixed, e.g. $(u_{1},u_{k})$, then we write $(u_{1},\ldots,u_{k})\sim{\cal N}_{\omega_{1},\ldots,\omega_{k-1}}|(u_{1},u_{k})$ to indicate that $(u_{1},\ldots,u_{k})$ is obtained by sampling from ${\cal N}_{\omega_{1},\ldots,\omega_{k-1}}$ conditioned on fixed values $(u_{1},u_{k})$. In that case $(u_{2},\ldots,u_{k-1})$ are random variables that depend on $(u_{1},u_{k})$.
3.2 Procedure ${\tt Matching}(u)$
In this section we describe a procedure that takes vector $u\in\mathbb{R}^{d}$ and either outputs a directed matching $M$ on nodes $V$ or terminates the oracle. In this procedure we choose constants $c^{\prime},\Delta,\sigma$ (to be specified later), and denote
$\pi$ $=$ $\frac{6\alpha}{c^{\prime}n\Delta}$ (10)
1
2compute $w_{x}=\langle x,u\rangle$ for each $x\in V$
3 sort $\{w_{x}\}_{x\in V}$, let $A,B$ be subsets of $V$ with $|A|=|B|=2c^{\prime}n$ containing nodes with the least and the greatest values of $w_{x}$, respectively
4 let $G^{\prime}$ be the graph obtained from $G$ by adding new vertices $s,t$ and edges $\{\{s,x\}\>:\>x\in A\}\cup\{\{y,t\}\>:\>y\in B\}$ of capacity $\pi$
5 compute maximum $s$-$t$ flow and the corresponding minimum $s$-$t$ cut in $G^{\prime}$
6 if capacity of the cut is less than $c^{\prime}n\pi=\tfrac{6\alpha}{\Delta}$ then return this cut and terminate the oracle
7 use flow decomposition to compute multicommodity flows $f_{p}$ and demands $d_{xy}$ (see text)
8 if flows $f_{p}$ satisfy condition (7) then return these flows and terminate the oracle
9 let $M_{\tt all}=\{(x,y)\in A\times B\>:\>d_{xy}>0,w_{y}-w_{x}\geq\sigma\}$ and $M_{\tt short}=\{(x,y)\in M_{\tt all}\>:\>||x-y||^{2}\leq\Delta\}$
pick maximal matching $M\subseteq M_{\tt short}$ and return $M$
Algorithm 1 ${\tt Matching}(u)$.
Let us elaborate line 6. Given flow $f^{\prime}$ in $G^{\prime}$, we compute its flow decomposition and remove flow cycles. Each path in this decomposition has the form $p^{\prime}=(s,p,t)$ where $p=(x,\ldots,y)$ with $x\in A$, $y\in B$. For each such $p$ we set $f_{p}=f^{\prime}_{p^{\prime}}$, and accordingly increase demand $d_{xy}$ by $f^{\prime}_{p^{\prime}}$. Note that we need to know only the endpoints of $p$, and not $p$ itself. This computation can be done in $O(m\log n)$ time using dynamic trees [ST83]. (The same subroutine was used in [She09]).
For the purpose of analysis we make the following assumption: if Algorithm 1 terminates at line 5 or 7 then it returns $\varnothing$ (the empty matching). Thus, we always have ${\tt Matching}(u)\subseteq V\times V$ and $|{\tt Matching}(u)|\leq|V|$.
Lemma 6.
(a) If the algorithm terminates at line 5 then the returned cut is $c^{\prime}$-balanced.
(b) There exist positive constants $c^{\prime},\sigma,\delta$ for which either (i) $\mathbb{E}_{u}|{\tt Matching}(u)|\geq\delta n$, or (ii) Algorithm 1 for $u\sim{\cal N}$ terminates at line 5 or 7 with probability at least $\Theta(1)$.
Proof.
(a)   We repeat the argument from [AK16]. Since the total flow is at most $\pi c^{\prime}n$, at least $c^{\prime}n$ newly added source edges are not saturated by the flow, and hence their endpoints are on the side of the source node in the cut obtained, which implies that that side has at least $c^{\prime}n$ nodes. Similarly, the other side of the cut also has at least $c^{\prime}n$ nodes, and thus the cut is $c^{\prime}$-balanced.
(b)   Assume that condition (ii) is false. By a standard argument, conditions (4a) and (4b) imply the following: there exist constants $c^{\prime}\in(0,c)$, $\beta>0$ and $\sigma>0$ such that with probability at least $\beta$ Algorithm 1 reaches line 9 and we have $w_{y}-w_{x}\geq\sigma$ for all $x\in A$, $y\in B$ (see [Kal07, Lemma 14]). Suppose that this event happens. We claim that in this case $|M|\geq\tfrac{1}{3}c^{\prime}n$. Indeed, suppose this is false. Let $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ be the sets of nodes involved in $M$ (with $|A^{\prime}|=|B^{\prime}|=|M|=k$). The total value of flow from $A$ to $B$ is at least $c^{\prime}\pi n$ (otherwise we would have terminated at line 5). The value of flow leaving $A^{\prime}$ is at most $|A^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Similarly, the value of flow entering $B^{\prime}$ is at most $|B^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Therefore, the value of flow from $A-A^{\prime}$ to $B-B^{\prime}$ is at least $c^{\prime}\pi n-2\cdot\tfrac{1}{3}c^{\prime}\pi n=\tfrac{1}{3}c^{\prime}\pi n$. For each edge $(x,y)\in M_{\tt all}$ with $x\in A-A^{\prime}$, $y\in B-B^{\prime}$ we have $||x-y||^{2}>\Delta$ (otherwise $M$ would not be a maximal matching in $M_{\tt short}$). Therefore,
$\sum_{p:p=(x,\ldots,y)}f_{p}||x-y||^{2}\geq\!\!\!\sum_{\begin{subarray}{c}p:p=(x,\ldots,y)\\ x\in A-A^{\prime},B\in B-B^{\prime}\end{subarray}}\!\!\!f_{p}||x-y||^{2}\geq\tfrac{1}{3}c^{\prime}\pi n\cdot\Delta=2\alpha$
But then the algorithm should have terminated at line 7 - a contradiction. ∎
In the remainder of the analysis we assume that case (i) holds in Lemma 6(b). (In the case of case (ii) procedures that we will describe will terminate the oracle at line 5 or 7 with probability $\Theta(1)$).
We assume that procedure ${\tt Matching}(\cdot)$ satisfies the following skew-symmetry condition: for any $u$, matching ${\tt Matching}(-u)$ is obtained from ${\tt Matching}(u)$ by reversing all edge orientations. (This can be easily enforced algorithmically).
3.3 Matching covers
We say that a generalized matching is a set $M$ of paths of the form $p=(p_{0},\ldots,p_{k})$ with $p_{0},\ldots,p_{k}\in V$ such that each node $x\in V$ has at most one incoming and at most one outgoing path. We say that path $q$ is violating if $q=(\ldots,p,\ldots)$ and path $p=(p_{0},\ldots,p_{\ell(p)})$ satisfies (9). We denote $M^{\tt violating}$ and $M^{\tt nonviolating}$ to be the set of violating and nonviolating paths in generalized matching $M$, respectively. Below we will only be interested in the endpoints of paths $p\in M$ and violating/nonviolating status of $p$. Thus, paths in $M$ will essentially be treated as edges with a Boolean flag. For generalized matchings $M_{1},M_{2}$ we define
$M_{1}\circ M_{2}=\{(p,x,q)\>:\>(p,x)\in M_{1},(x,q)\in M_{2}\}$
where $p,q$ are paths and $x$ is a node. Clearly, $M_{1}\circ M_{2}$ is also a generalized matching. For generalized matching $M$, vector $u\in\mathbb{R}^{d}$ and value $\sigma\in\mathbb{R}$ we define
${\tt Truncate}_{\sigma}(M;u)$ $=M^{\tt violating}\;\;\cup\;\;\{(x,\ldots,y)\in M^{\tt nonviolating}\>:\>\langle y-x,u\rangle\geq\sigma\}$
We will consider algorithms for constructing generalized matchings that have the following form: given vector $u\in\mathbb{R}^{d}$, sample vectors $(u_{1},\ldots,u_{k})$ according to some distribution that depends on $u$, and return ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{k})$. Any such algorithm specifies a matching cover as defined below.
Definition 7.
A generalized matching cover (or just matching cover) is function ${\cal M}$ that maps vector $u\in\mathbb{R}^{d}$ to a distribution over generalized matchings. It is called skew-symmetric if sampling $M\sim{\cal M}(-u)$ and then reversing all paths in $M$ produces the same distribution as sampling $M\sim{\cal M}(u)$. We define ${\tt size}({\cal M})=\tfrac{1}{n}\mathbb{E}_{M\sim{\cal M}({\cal N})}[|M|]$ where ${\cal M}({\cal N})$ denotes the distribution $u\sim{\cal N},M\sim{\cal M}(u)$. We say that ${\cal M}$ is $\sigma$-stretched (resp. $L$-long) if $\langle y-x,u\rangle\geq\sigma$ (resp. $||y-x||^{2}\leq L$) for any $u\in\mathbb{R}^{d}$ and any nonviolating $(x,\ldots,y)\in{\tt supp}({\cal M}(u))$. ${\cal M}$ is $k$-hop if any path $p\in M\in{\tt supp}({\cal M}({\cal N}))$ has the form $p=(p_{0},p_{1},\ldots,p_{k})$ where $||p_{i}-p_{i-1}||^{2}\leq\Delta$ for all $i\in[k]$. We write ${\cal M}\subseteq{\cal M}^{\prime}$ for (coupled) matching covers ${\cal M},{\cal M}^{\prime}$ if $M\subseteq M^{\prime}$ for any $u$ and $M\sim{\cal M}(u)$, $M^{\prime}\sim{\cal M}(u^{\prime})$.
If ${\cal M}$ is a matching cover then we define matching cover ${\cal M}_{[\sigma]}$ as follows: given vector $u$, sample $M\sim{\cal M}(u)$ and let ${\tt Truncate}_{\sigma}(M;u)$ be the output of ${\cal M}_{[\sigma]}(u)$. Clearly, ${\cal M}_{[\sigma]}$ is $\sigma$-stretched.
The following lemma shows how matching covers can be used. We say that direction $u\in\mathbb{R}^{d}$ is regular if $\langle y-x,u\rangle<\sqrt{6\ln n}\cdot||y-x||$ for all distinct $x,y\in V$.
Lemma 8.
Let ${\cal M}$ be a $k$-hop matching cover.
(a) ${\cal M}$ is $((k+1)\Delta)$-long.
(b) If ${\cal M}$ is $\sqrt{6(k+1)\Delta\ln n}$-stretched then $M^{\tt nonviolating}\!=\!\varnothing$ for any regular $u$ and $M\!\in\!{\tt supp}({\cal M}(u))$.
(c) ${\tt Pr}_{u}[u\mbox{ is regular\,}]\geq 1-\tfrac{1}{n}$.
Proof.
(a)   By definitions, any nonviolating path $p=(p_{0},p_{1},\ldots,p_{k})\in M\in{\tt supp}({\cal M}({\cal N}))$ satisfies $||p_{k}-p_{0}||^{2}\leq\Delta+\sum_{i=1}^{k}||p_{i}-p_{i-1}||^{2}\leq\Delta+\sum_{i=1}^{k}\Delta=(k+1)\Delta$.
(b)   Suppose there exists nonviolating path $(x,\ldots,y)\in M\in{\tt supp}({\cal M}(u))$. Elements $x,y$ must be distinct. Part (a) gives $||y-x||\leq\sqrt{(k+1)\Delta}$. Regularity of $u$ thus implies that $\langle y-x,u\rangle<\sqrt{6\ln n}\cdot\sqrt{(k+1)\Delta}$ a contradiction.
(c)   Consider distinct $x,y\in V$. The quantity $\langle y-x,u\rangle$ is normal with zero mean and variance $||y-x||^{2}$ under $u\sim{\cal N}$. Thus, ${\tt Pr}_{u}[\langle y-x,u\rangle\geq c||y-x||]={\tt Pr}[{\cal N}(0,||y-x||)\geq c||y-x||]={\tt Pr}[{\cal N}(0,1)\geq c]\leq e^{-c^{2}/2}=\tfrac{1}{n^{3}}$ for $c=\sqrt{6\ln n}$. There are at most $n^{2}$ distinct pairs $x,y\in V$, so the union bound gives the claim. ∎
3.4 Chaining algorithms
By construction, ${\tt Matching}$ is a (deterministic) 1-hop $\sigma$-stretched skew-symmetric matching cover. Next, we discuss how to “chain together” matchings returned by ${\tt Matching}(\cdot)$ to obtain a matching cover with a large stretch, as required by Lemma 8(b).
Sherman’s algorithm   First, we review Sherman’s algorithm [She09]. In addition to vector $u$, it takes a sequence $b=(b_{1},\ldots,b_{K})\in\{0,1\}^{K}$ as an input.
1 let $k$ be the number of $1$’s in $b$ and $k^{\prime}$ be the number of 00’s in $b$, so that $k+k^{\prime}=K$
2 sample $(u_{1},\ldots,u_{k})\sim{\cal N}_{\omega}^{k}|u_{1}$ and $(u^{\prime}_{1},\ldots,u^{\prime}_{k^{\prime}})\sim{\cal N}_{0}^{k^{\prime}}$ independently where $\omega=1-1/k$
3 let $(\bar{u}_{1},\ldots,\bar{u}_{K})$ be the sequence obtained by merging sequences $(u_{1},\ldots,u_{k})$ and $(u^{\prime}_{1},\ldots,u^{\prime}_{k^{\prime}})$ at positions specified by $b$ in a natural way
return ${\tt Matching}(\bar{u}_{1})\circ\ldots\circ{\tt Matching}(\bar{u}_{K})$
Algorithm 2 ${\tt SamplePaths}^{b}(u_{1})$.
Theorem 9.
For any $\delta>0$ and $\sigma>0$ there exist positive constants $c_{1},c_{2},c_{3}$ with the following property. Suppose that ${\tt Matching}$ is a 1-hop $\sigma$-stretched skew-symmetric matching cover with ${\tt size}({\tt Matching})\geq\delta$, and $K\Delta\leq c_{1}$. Then there exists vector $b\in\{0,1\}^{K}$ for which ${\tt size}\left({\tt SamplePaths}^{b}_{[c_{2}K]}\right)\geq e^{-c_{3}K^{2}}$.
In practice vector $b$ is not known, however it can be sampled uniformly at random which decreases the expectation by a factor of $2^{-K}$, which does not change the bound $e^{-\Theta(K^{2})}$ in Theorem 9. Note that [She09] does not use the notion of “violating paths”, and accordingly Theorem 9 is formulated slightly differently in [She09]. However, the proof in [She09] can be easily adapted to yield Theorem 9333 We believe that [She09] has a numerical mistake. Namely, consider value $\delta\in(0,1)$ and weighted graph $(V,E,w)$ with $n=|V|$ nodes and non-negative weights $w$ such that $w(x,A)=w(A,x)\leq 1$ for all $x\in V$ and $A\subseteq V$ (where $w(X,Y)=\sum_{(x,y)\in X\times Y}w(x,y)$). [She09, proof of Lemma 4.8] essentially makes the following claim: For any subset $B\subseteq V$ with $w(V,B)\geq\delta|B|$ there exists $A\subseteq V$ such that either (i) $|A|\geq\delta|B|$ and $w(x,B)\geq\delta^{3}$ for any $x\in A$, or (ii) $|A|\geq\tfrac{1}{\delta}|B|$ and $w(x,B)\geq\tfrac{\delta^{2}}{n}|B|$ for any $x\in A$. A counterexample can be constructed as follows. Assume that $\delta|B|$ and $\tfrac{1}{\delta^{4}}$ are integers. For $\delta|B|-1$ nodes $x$ set $w(x,B)=1$, for $\tfrac{1}{\delta^{4}}$ nodes $x$ set $w(x,B)=\delta^{4}$, and for remaining nodes set $w(x,B)=0$. Then the claim above is false if $\delta|B|-1+\tfrac{1}{\delta^{4}}<\tfrac{1}{\delta}|B|$. The statement can be corrected as follows: For any $\lambda\leq\delta$ and any subset $B\subseteq V$ with $w(V,B)\geq\delta|B|$ there exists $A\subseteq V$ such that either (i) $|A|\geq\tfrac{\delta}{3}|B|$ and $w(x,B)\geq\lambda$ for any $x\in A$, or (ii) $|A|\geq\tfrac{\delta}{3\lambda}|B|$ and $w(x,B)\geq\tfrac{\delta}{3n}|B|$ for any $x\in A$. (If both conditions are false then $w(V,B)<\tfrac{\delta}{3}|B|\cdot 1+\tfrac{\delta}{3\lambda}|B|\cdot\lambda+n\cdot\tfrac{\delta}{3n}|B|=\delta|B|$, which is impossible since $w(V,y)\geq\delta$ for each $y\in B$). Then the proof in [She09] still works, but with different constants.
Robustness to deletions and parallelization   Note that Theorem 9 can also be applied to any matching cover ${\tt Matching}^{\prime}\subseteq{\tt Matching}$ satisfying ${\tt size}({\tt Matching}^{\prime})\geq\delta^{\prime}$ for some positive constant $\delta^{\prime}<\delta$. Thus, we can adversarially “knock out” some edges from ${\tt Matching}$ and still get useful bounds on the size of the output. This is a key proof technique in this paper; to our knowledge, it has not been exploited before. Our first use of this technique is for parallelization. We need to show that running Algorithm 2 multiple times independently produces many disjoint paths. Roughly speaking, our argument is as follows. Consider the $i$-th run, and assume that $|M_{\tt prev}|$ is small where $M_{\tt prev}$ is the union of paths computed in the first $i-1$ runs. Let us apply Theorem 9 to the matching cover obtained from ${\tt Matching}$ by knocking out edges incident to nodes in $M_{\tt prev}$. It yields that the $i$-th run produces many paths that are node-disjoint from $M_{\tt prev}$, as desired. We refer to Section 3.5 for further details.
New chaining algorithm   Next, we discuss how a similar proof technique (combined with additional ideas) can be used to simplify Sherman’s chaining algorithm. We will show that Theorem 9 still holds if we consider vectors $b$ of the form $b=(1,\ldots,1)$. The algorithm thus becomes as follows.
sample $(u_{1},\ldots,u_{K})\sim{\cal N}_{\omega}^{K}|u_{1}$ where $\omega=1-1/K$
  // equivalently, sample $u_{2}\!\sim_{\omega}\!u_{1}$, $u_{3}\!\sim_{\omega}\!u_{2},~{}\ldots,~{}u_{K}\!\sim_{\omega}\!u_{K-1}$
return ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{K})$
Algorithm 3 ${\tt SamplePaths}^{K}(u_{1})$.
Theorem 10.
For any $\delta>0$ and $\sigma>0$ there exist positive constants $c_{1},c_{2},c_{3}$ with the following property. Suppose that ${\tt Matching}$ is a 1-hop $\sigma$-stretched skew-symmetric matching cover with ${\tt size}({\tt Matching})\geq\delta$, $K\Delta\leq c_{1}$, and $K=2^{r}$ for some integer $r$. Then ${\tt size}\left({\tt SamplePaths}^{K}_{[c_{2}K]}\right)\geq e^{-c_{3}K^{2}}$.
We prove this theorem in Section 4. Our proof technique is very different from [She09], and relies on two key ideas:
  1. 1.
    We use induction on $k=2^{0},2^{1},2^{2},\ldots,K$ to show that ${\tt SamplePaths}^{k}_{[\sigma_{k}]}$ has a sufficiently large size assuming that ${\tt size}({\tt Matching})\geq\delta_{k}$, for some sequences $\delta_{1}<\delta_{2}<\delta_{4}<\ldots<\delta_{K}=\delta$ and $\sigma_{1}<\sigma_{2}<\sigma_{4}<\ldots<\sigma_{K}=\Theta(K)$. To prove the claim for $2k$, we show that for each node $x$, function $\mu_{x}(u)$ is “sufficiently spread” where $\mu_{x}(u)$ is the expected out-degree of $x$ in $M\sim{\tt SamplePaths}^{k}_{[\sigma_{k}]}(u)$. In order to do this, we “knock out” edges $(x,y)$ from ${\tt Matching}(u)$ for a $\Theta(\delta_{2k}-\delta_{k})$ fraction of vectors $u$ with the largest value of $\mu_{x}(u)$, and then use the induction hypothesis for the smaller matching cover of size $\delta_{k}$. We conclude that $\mu_{x}(u)$ is sufficiently large for $\Theta(\delta_{2k}-\delta_{k})$ fraction of $u$’s. We then use Theorem 5 and skew-symmetry to argue that chaining ${\tt SamplePaths}^{k}_{[\sigma_{k}]}$ with itself gives a matching cover of large size.
  2. 2.
    We work with “extended matching covers” instead of matching covers. These are functions ${\cal M}$ that take a pair of vectors $(u_{1},u_{k})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ as input, sample $(u_{1},\ldots,u_{k})\sim{\cal N}^{k}_{\omega}$ conditioned on fixed $u_{1},u_{k}$, and return (a subset of) ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{k})$. This guarantees that if $(x,y)$ is removed from ${\tt Matching}(u_{1})$ then $x$ has no outgoing edge in ${\cal M}(u_{1},u_{k})$ (which is needed by argument above).
Our proof appears to be more compact than Sherman’s proof, and also does not rely on case analysis. Accordingly, we believe that our technique should give smaller constants in the $O(\cdot)$ notation (although neither proof explicitly optimizes these constants).
Next, we discuss implications of Theorem 10. Below we denote $\llceil z\rrceil=2^{\lceil\log_{2}z\rceil}$ to be the smallest $K=2^{r}$, $r\in\mathbb{Z}_{\geq 0}$ satisfying $K\geq z$.
Corollary 11.
Suppose that ${\tt Matching}$ is a 1-hop $\sigma$-stretched skew-symmetric matching cover with ${\tt size}({\tt Matching})\geq\delta$. Define $A=\tfrac{12}{c_{2}^{2}}$, $B=\frac{1}{2A\sqrt{c_{3}}}$ and $\varepsilon^{\max}=\frac{c_{1}}{2AB^{2}}$. Suppose that $\Delta=B\sqrt{\frac{\varepsilon}{\ln n}}$ where $\varepsilon\in(0,\varepsilon^{\max}]$ and $A\Delta\ln n\geq 1$. If $K=\llceil A\Delta\ln n\rrceil$ then
$\mathbb{E}_{M\sim{\tt SamplePaths}^{K}({\cal N})}[|M^{\tt violating}|]\geq n^{1-\varepsilon}-1$
Proof.
Note that $K\in[A\Delta\ln n,2A\Delta\ln n]$, and $K\Delta\leq 2A\Delta^{2}\ln n=2AB^{2}\varepsilon\leq 2AB^{2}\varepsilon^{\max}=c_{1}$. Denote ${\cal M}={\tt SamplePaths}^{K}_{[c_{2}K]}$. Theorem 10 gives ${\tt size}({\cal M})\geq e^{-c_{3}K^{2}}\geq e^{-c_{3}(2A\Delta\ln n)^{2}}=n^{-\varepsilon}$. We also have $\frac{6(K+1)\Delta\ln n}{(c_{2}K)^{2}}\leq\frac{12\Delta\ln n}{c_{2}^{2}K}\leq\frac{12\Delta\ln n}{c_{2}^{2}A\Delta\ln n}=1$ and so the precondition of Lemma 8(b) holds for ${\cal M}$.
Let $x=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M|]=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M^{\tt violating}|]$ (by Lemma 8(b)), $y=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is not regular},M\sim{\cal M}(u)}[|M|]\leq n$ and $p={\tt Pr}_{u}[u\mbox{ is not regular}]\leq\tfrac{1}{n}$ (by Lemma 8(c)). We have $n^{1-\varepsilon}\leq\mathbb{E}_{M\sim{\cal M}({\cal N})}[|M|]=(1-p)x+py\leq(1-p)x+\tfrac{1}{n}\cdot n$ and so $(1-p)x\geq n^{1-\varepsilon}-1$. Therefore, $\mathbb{E}_{M\sim{\tt SamplePaths}^{K}({\cal N})}[|M^{\tt violating}|]\geq(1-p)x\geq n^{1-\varepsilon}-1$.
3.5 Final algorithm
In the algorithm below we use the following notation: $V(p)\subseteq V$ is the set of nodes through which path $p$ passes, and $V(M)=\bigcup_{p\in M}V(p)$. Furthermore, if $p$ is violating then $p^{\tt violating}$ is a subpath of $p$ satisfying (8). Note that lines 1-3 compute sets of paths $\tilde{M}_{1},\ldots,\tilde{M}_{N}$, which are then combined into a single set $M\subseteq\bigcup_{i}\tilde{M}_{i}$ using one of the two options. Option 1 will be mainly used for the analysis, while option 2 will be used for an efficient parallel implementation.
1 for $i=1,\ldots,N$ do
2       sample $u\sim{\cal N}$, call $M_{i}={\tt SamplePaths}^{K}(u)$
3       let $\tilde{M}_{i}=\{p^{\tt violating}\>:\>p\in M_{i}^{\tt violating}\}$
4option 1: set $M\!:=\!\varnothing$, then for $i\!=\!1,\ldots,N$ update $M\!:=\!M\cup\{p\in\tilde{M}_{i}\>:\>V(p)\cap V(M)=\varnothing\}$
[…]
return $M$
Algorithm 4 Computing violating paths.
Theorem 12.
Suppose that ${\tt Matching}$ is a 1-hop $\sigma$-stretched skew-symmetric matching cover with ${\tt size}({\tt Matching})\geq\delta$, and let $A,B,\varepsilon^{\max}$ be the constants defined on Corollary 11 for value $\delta/2$. Suppose that $\Delta=B\sqrt{\frac{\varepsilon}{\ln n}}$ where $\varepsilon\in(0,\varepsilon^{\max}]$ and $A\Delta\ln n\geq 1$. If $K=\llceil A\Delta\ln n\rrceil$ and $N\geq\tfrac{\delta n^{\varepsilon}}{4K(1-n^{\varepsilon-1})}$ then $\mathbb{E}[|M|]\geq\tfrac{\delta n}{8K}$ where $M$ is the output of Algorithm 4 with option 1.
Proof.
Let ${\cal E}_{i}$ be the event that set $M$ at the beginning of iteration $i$ satisfies $|M|\leq a:=\tfrac{\delta n}{4K}$, and let $\gamma_{i}={\tt Pr}[{\cal E}_{i}]$. Let $x_{i}$ be the expected number of paths that have been added to $M$ at iteration $i$, so that $\mathbb{E}[|M|]=x_{1}+\ldots+x_{N}$ for the final set $M$. Let $y_{i}$ be the expected number of paths that have been added to $M$ at iteration $i$ conditioned on event ${\cal E}_{i}$. Clearly, we have $x_{i}\geq\gamma_{i}y_{i}\geq\gamma y_{i}$ where $\gamma:=\gamma_{N}$.
Next, we bound $y_{i}$. Let $M$ be the set at the beginning of iteration $i$, and suppose that ${\cal E}_{i}$ holds, i.e. $|M|\leq\tfrac{\delta n}{4K}$. Denote $U=V(M)$, then $|U|\leq K|M|\leq\tfrac{\delta n}{4}$. Let ${\tt Matching}^{\prime}\subseteq{\tt Matching}$ be the (skew-symmetric) matching cover obtained by removing from ${\tt Matching}(u)$ edges $(x,y)$ and $(y,x)$ with $x\in U$ (for all $u\in\mathbb{R}^{d}$). Clearly, we have ${\tt size}({\tt Matching}^{\prime})\geq\delta-\tfrac{1}{2}\delta=\tfrac{1}{2}\delta$. Let ${\tt SamplePaths}^{\prime K}$ be the matching cover given by Algorithm 3 where ${\tt Matching}$ is replaced by ${\tt Matching}^{\prime}$. Clearly, we have ${\tt SamplePaths}^{\prime K}\subseteq{\tt SamplePaths}^{K}$. By Corollary 11 applied to ${\tt Matching}^{\prime}$, ${\tt SamplePaths}^{\prime K}({\cal N})$ produces at least $n^{1-\varepsilon}-1$ violating paths in expectation. By construction, all these paths $p$ satisfy $V(M)\cap V(p)=\varnothing$, and therefore $y_{i}\geq n^{1-\varepsilon}-1$.
We showed that $\mathbb{E}[|M|]\geq\sum_{i=1}^{N}\gamma(n^{1-\varepsilon}-1)=\gamma b$ for the final set $M$ where $b:=N(n^{1-\varepsilon}-1)$. We also have $\mathbb{E}[|M|]\geq(1-\gamma)a$, and hence $\mathbb{E}[|M|]\geq\min_{\gamma\in[0,1]}\max\{\gamma b,(1-\gamma)a\}=\tfrac{ab}{a+b}$ (the minimum is attained at $\gamma=a/(a+b)$). By assumption, we have $b\geq a$, and so $\mathbb{E}[|M|]\geq\tfrac{1}{2}a$. ∎
Next, we analyze Algorithm 4 with option 2. Recall that “parallel runtime” is the maximum runtime over all processors.
Lemma 13.
(a) Let $M$ and $M^{\prime}$ be the outputs of Algorithm 4 with options 1 and 2, respectively (for a given run of the loop in lines 1-3). Then $|M^{\prime}|\geq|M|/K^{3}$.
(b) Algorithm 4 with option 2 can be implemented on $N$ processors with parallel runtime $O(KT_{\tt maxflow}+n(Kd+K\log^{2}n+\log N)))$ (in any version of PRAM).
Proof.
(a)   By the construction of $M$, each node $v\in V(M^{\ast})$ is contained in at most $K$ paths $p\in M$ (all of them belong to some $\tilde{M}_{i}$ for fixed $i$). Therefore, $|V(M^{\ast})|\geq|M|/K$. We also have $|V(M^{\prime})|\geq|V(M^{\ast})|/K$ and $|M^{\prime}|\geq|V(M^{\prime})|/K$, since $|V(p)|\leq K$ for any $p\in M^{\ast}$. Putting these inequalities together gives the claim.
(b)   Clearly, sets $\tilde{M}_{i}$ for $i\in[N]$ can be computed in parallel on $N$ processors in time $O(K(T_{\tt maxflow}+nd+n\log^{2}n)))$ per processor (since computing dot products $\langle x,u\rangle$ in Algorithm 1 takes time $O(Knd)$, and flow decompositions take time $O(Km\log n)=O(Kn\log^{2}n)$). We can assume that after computing $\tilde{M}_{i}$, processor $i$ computes a maximal set of paths $\tilde{M}_{i}^{\prime}\subseteq\tilde{M}_{i}$ that are pairwise node-disjoint, and updates $\tilde{M}_{i}:=\tilde{M}_{i}^{\prime}$. Clearly, this step does not affect the output of line 5.
It remains to discuss how to implement line 5 (computing a maximal set of paths $M$ in $M^{\ast}$ which are pairwise node-disjoint). For $N=2$ this can be easily done in $O(n)$ time: processor 1 sends $\tilde{M}_{1}$ to processor 2, and processor 2 computes the answer. The general case can be reduced to the case above using a divide-and-conquere strategy with a computation tree which is a binary tree whose leaves are the $N$ processors. The depth of this tree is $O(\log N)$, and hence the parallel runtime of this procedure is $O(n\log N)$. ∎
Note in Algorithm 4 step 3 can be run in parallel on $N$ processors; after all of them finish, we can run the rest of algorithm on a single machine. By putting everything together, we obtain
Theorem 14.
There exists an algorithm for Balanced Separator that given $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$, produces $O(\sqrt{(\log n)/\varepsilon})$-pseudoapproximation w.h.p.. Its expected parallel runtime is $O((\log^{O(1)}n)T_{\tt maxflow})$ on $O(n^{\varepsilon})$ processors (in any version of PRAM).
Proof.
Let $\delta,\sigma$ be as in Lemma 6. Set parameters as in Theorem 12, and require additionally that $\varepsilon\leq\tfrac{1}{2}$. Condition $A\Delta\ln n\geq 1$ means that this can be done for $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$. By Theorem 12 and Lemma 13, the output of Algorithm 4 satisfies $\mathbb{E}[|M|]\geq\frac{\delta n}{8K^{1+h}}$ where $h=0$ if option 1 is used, and $h=3$ if option 2 is used.
To implement the oracle, run Algorithm 4 until either procedure ${\tt Matching}(\cdot)$ (Alg. 1) terminates at lines 5 or 7, or until we find a set $M$ of violating paths with $|M|\geq\tfrac{\delta n}{16K^{1+h}}$. In the latter case use Lemma 3 to set the variables. Since we always have $|M|\leq n$, the expected number of runs will be $O(n/\tfrac{n}{K^{1+h}})=O(K^{1+h})$.
If the oracle terminates at line 5 of Alg. 1, then it returns a $c^{\prime}$-balanced cut of cost at most $\tfrac{6\alpha}{\Delta}=O\left(\alpha\sqrt{\tfrac{\log n}{\varepsilon}}\right)$. Otherwise it returns valid variables $f_{p},F$. If the oracle terminates at line 7 then its width is $\rho=O(\tfrac{\alpha}{n}+\pi)=O(\tfrac{\alpha}{n\Delta})$. Now suppose that the oracle finds set $M$ of violating paths with $|M|\geq\tfrac{\delta n}{16K^{1+h}}$. Clearly, degrees $\pi_{F},\pi_{D}$ in Lemma 3 satisfy $\pi_{F}=O(K)$ and $\pi_{D}=O(1)$, so the oracle’s width in this case is $\rho=O(\tfrac{\alpha}{n}+\tfrac{\alpha K^{2+h}}{n\Delta})$. In both cases we have $\rho=O(\tfrac{\alpha K^{2+h}}{n\Delta})$. Thus, the number of calls to ${\tt Oracle}$ for a fixed value of $\alpha$ is $T=O(\tfrac{\rho^{2}n^{2}\log n}{\alpha^{2}})=O(\tfrac{K^{4+2h}\log n}{\Delta^{2}})$.
Next, we bound the complexity of computing approximations $\tilde{v}_{1},\ldots,\tilde{v}_{n_{\tt orig}}$ for fixed $\alpha$ as described in Theorem 21 in Section A.1. (Here we assume familiarity with Appendix A). We have $\tau=\Theta(1)$ and $\gamma=\Theta\{\min\{\tfrac{\alpha}{n\pi},\tfrac{\Delta}{K}\})=\Theta(\tfrac{\Delta}{K})$, thus we need to use dimension $d=\Theta(\tfrac{\log n}{\gamma^{2}})=\Theta(\tfrac{K^{2}\log n}{\Delta^{2}})$. We need to compute $Tkd$ matrix-vector products of form $A\cdot u$ where $k=O(\max\{(\tfrac{\rho n\log n}{\alpha})^{2},\log n\})=O(\tfrac{K^{4}\log^{2}n}{\Delta^{2}})$. Each matrix $A$ has the form $\sum_{r=1}^{O(T)}N^{(r)}$, and each $N^{(r)}$ can be represented as a sum of $O(K)$ “easy” matrices (e.g. corresponding to matchings) for which the multiplication with a vector takes $O(n)$ time. To summarize, the overall complexity of computing approximations $\tilde{v}_{1},\ldots,\tilde{v}_{n_{\tt orig}}$ is $Tkd\cdot TKn=O(\tfrac{K^{15+4h}\log^{5}n}{\Delta^{8}}n)$, which is $O(n\log^{O(1)}n)$ since $\Delta=\Theta\left(\sqrt{\tfrac{\varepsilon}{\log n}}\right)\in\left[\Theta(\tfrac{1}{\log n}),\Theta(\tfrac{1}{\sqrt{\log n}})\right]$ and $K=\Theta(\Delta\log n)=O(\sqrt{\log n})$. These computations can be done on a single processor; their runtime is subsumed by the claimed $O((\log^{O(1)}n)T_{\tt maxflow})$ bound.
We saw that $d=O(\log^{O(1)}n)$. From Lemma 13 we can now conclude that the algorithm’s expected parallel runtime is $T\cdot O(KT_{\tt maxflow}+n(Kd+K\log^{2}n+\log N)))=O((\log^{O(1)}n)T_{\tt maxflow})$ on $N=\Theta(n^{\varepsilon}/K)$ processors, assuming that option 2 is used. Note that the output of the overall algorithm is correct w.h.p. since vectors $\tilde{v}_{1},\ldots,\tilde{v}_{n_{\tt orig}}$ approximate original vectors only w.h.p. (see Theorem 21). ∎
4 Proof of Theorem 10
We will need the following definition.
Definition 15.
A $k$-hop extended matching cover is function ${\cal M}$ that maps vectors $u,u^{\prime}\in\mathbb{R}^{d}$ to a distribution over generalized matchings. It is skew-symmetric if sampling $M\sim{\cal M}(-u^{\prime},-u)$ and then reversing all paths in $M$ produces the same distribution as sampling $M\sim{\cal M}(u,u^{\prime})$. We define ${\tt size}_{\omega}({\cal M})=\tfrac{1}{n}\mathbb{E}_{M\sim{\cal M}({\cal N}_{\omega})}[|M|]$ where ${\cal M}({\cal N}_{\omega})$ denotes the distribution $(u,u^{\prime})\sim{\cal N}_{\omega},M\sim{\cal M}(u,u^{\prime})$. We say that ${\cal M}$ is $\sigma$-stretched (resp. $L$-long) if $\min\{\langle y-x,u\rangle,\langle y-x,u^{\prime}\rangle\}\geq\sigma$ (resp. $||y-x||^{2}\leq L$) for any $u,u^{\prime}\in\mathbb{R}^{d}$ and any nonviolating $(x,\ldots,y)\in{\tt supp}({\cal M}(u,u^{\prime}))$. We write ${\cal M}\subseteq\tilde{\cal M}$ for (coupled) extended matching covers ${\cal M},\tilde{\cal M}$ if $M\subseteq\tilde{M}$ for any $u,u^{\prime}$ and $M\sim{\cal M}(u,u^{\prime})$, $\tilde{M}\sim\tilde{\cal M}(u,u^{\prime})$.
For an extended matching cover ${\cal M}$ we can define matching cover ${\cal M}_{\omega}$ as follows: given vector $u$, sample $u^{\prime}\sim_{\omega}u$, $M\sim{\cal M}(u,u^{\prime})$ and let $M$ be the output of ${\cal M}(u)$. Clearly, if ${\cal M}$ is $\sigma$-stretched then ${\cal M}_{\omega}$ is also $\sigma$-stretched, and ${\tt size}_{\omega}({\cal M})={\tt size}({\cal M}_{\omega})$.
We now proceed with the proof of Theorem 10. Define ${\cal K}=\{2^{0},2^{1},2^{2},\ldots,K\}$, and let $\dot{\cal K}={\cal K}-\{K\}$. (Recall that $K$ has the form $K=2^{r}$ for some integer $r$). Let us choose positive numbers $\beta_{k}$ for $k\in\dot{\cal K}$ (to be specified later), and define numbers $\{\sigma_{k}\}_{k\in{\cal K}}$ via the following recursion:
$\sigma_{1}$ $=$ $\sigma$ (11a)
$\sigma_{2k}$ $=$ $(1+\omega^{k})\sigma_{k}-\beta_{k}\qquad\forall k\in\dot{\cal K}$ (11b)
For a matching cover ${\cal M}\subseteq{\tt Matching}$ and integer $k\in{\cal K}$ we define extended matching cover ${\cal M}^{k}$ as follows:
  • given vectors $(u_{1},u_{k})$, sample $(u_{1},u_{2},\ldots,u_{k})\sim{\cal N}_{\omega}^{k}|(u_{1},u_{k})$, compute $M={\cal M}(u_{1})\circ\ldots\circ{\cal M}(u_{k})$ and let ${\tt Truncate}_{\sigma_{k}}(M;u_{1})\cap{\tt Truncate}_{\sigma_{k}}(M;u_{k})$ be the output of ${\cal M}^{k}(u_{1},u_{k})$.
By definition, ${\cal M}^{k}$ is $\sigma_{k}$-stretched. It can be seen that $({\cal M}^{K})_{\omega^{K-1}}\subseteq{\tt SamplePaths}^{K}$ for any ${\cal M}\subseteq{\tt Matching}$. Our goal will be to analyze ${\tt size}_{\omega^{K-1}}({\cal M}^{K})={\tt size}(({\cal M}^{K})_{\omega^{k-1}})$ for ${\cal M}={\tt Matching}$.
Theorem 16.
Choose an increasing sequence of numbers $\{\delta_{k}\}_{k\in{\cal K}}$ in $(0,1)$, and define sequence $\{\lambda_{k}\}_{k\in{\cal K}}$ via the following recursions:
[…]
Suppose that
$\exp\left(-\frac{\beta_{k}^{2}}{2(k+1)(1-\omega^{k})\Delta}\right)\leq\tfrac{1}{2}\lambda_{2k}\qquad\forall k\in\dot{\cal K}$ (13)
Then for any $k\in{\cal K}$ and any skew-symmetric matching cover ${\cal M}\subseteq{\tt Matching}$ with ${\tt size}({\cal M})\geq\delta_{k}$ there holds ${\tt size}_{\omega^{k-1}}({\cal M}^{k})\geq\lambda_{k}$.
By plugging appropriate sequences $\{\delta_{k}\}$ and $\{\beta_{k}\}$ we can derive Theorem 10 as follows.
Lemma 17.
Define $\delta_{k}=(1-\tfrac{1}{2k})\delta$ and $\beta_{k}=2\phi\sigma k\sqrt{1-\omega^{k}}$ where $\phi=\tfrac{1}{\sigma}\left(K\Delta\cdot 5\ln\tfrac{16}{\delta}\right)^{1/2}$. Then (13) holds, and
$\lambda_{k}$$\geq$ $\left(\frac{16}{\delta}\right)^{-2Kk}$ (14)
Furthermore, if $\phi\leq\tfrac{1}{8}$ (or equivalently $K\Delta<\frac{\sigma^{2}}{320\ln\tfrac{16}{\delta}}$ ) then
$\sigma_{K}$$\geq$ $\tfrac{1}{16}\,\sigma\,K$ (15)
Proof.
Recall that $\omega=1-1/K$. Let us set $\delta_{k}=(1-\tfrac{1}{2k})\delta$, then $\delta_{1}=\tfrac{1}{2}\delta$ and $\theta_{j}=\tfrac{1}{2}\left(\tfrac{\delta}{4j}\right)^{2K}=\tfrac{C}{j^{2K}}$ for $j\geq 2$ where $C=\tfrac{1}{2}\left(\tfrac{\delta}{4}\right)^{2K}$. By expanding expressions (12) we obtain
$\lambda_{k}$ $=$ $\delta_{1}^{k}\prod_{i\in{\cal K}\cap[k/2]}\theta^{i}_{k/i}\;\;=\;\;\left(\frac{\delta}{2}\right)^{k}\cdot\left(\prod_{i\in{\cal K}\cap[k/2]}C^{i}\right)\cdot\left(\prod_{i\in{\cal K}\cap[k/2]}(i/k)^{i}\right)^{2K}$
>$\stackrel{{\scriptstyle\mbox{\tiny(1)}}}{{>}}$ $\left(\frac{\delta}{2}\right)^{k}\cdot C^{k-1}\cdot\left(2^{-2k}\right)^{2K}\;\;\stackrel{{\scriptstyle\mbox{\tiny(2)}}}{{>}}\;\;(2C)^{k}\cdot 2^{-4Kk}\;\;=\;\;\left(\frac{16}{\delta}\right)^{-2Kk}\qquad\quad\forall k\in{\cal K}-\{1\}$
where in (2) we used the fact that $\left(\tfrac{\delta}{2}\right)^{k}>2^{k}\cdot C$, and in (1) we used the fact that for $k=2^{r}\in{\cal K}$ we have
$\log_{2}\prod_{i\in{\cal K}\cap[k/2]}(k/i)^{i}=\sum_{i\in{\cal K}\cap[k/2]}i\log(k/i)=\sum_{j\in{\cal K}\cap[2,k]}(k/j)\log j=\sum_{i=1}^{r}2^{r-i}\cdot i<2^{r}\cdot 2=2k$
since $\sum_{i=1}^{\infty}i2^{-i}=\tfrac{1}{2}f(\tfrac{1}{2})=2$ where $f(z)=\tfrac{1}{(1-z)^{2}}=(\tfrac{1}{1-z})^{\prime}=\sum_{i=1}^{\infty}iz^{i-1}$ for $|z|<1$.
Next, we show (13). We have
$\beta_{k}^{2}=K\Delta\cdot 20\ln\tfrac{16}{\delta}\cdot k^{2}(1-\omega^{k})\geq K\Delta\cdot 10\ln\tfrac{16}{\delta}\cdot k(k+1)(1-\omega^{k})$
and so
$\exp\left(-\frac{\beta_{k}^{2}}{2(k+1)(1-\omega^{k})\Delta}\right)\leq\exp\left(-Kk\cdot 5\ln\tfrac{16}{\delta}\right)=\left(\tfrac{16}{\delta}\right)^{-5Kk}\leq\tfrac{1}{2}\left(\tfrac{16}{\delta}\right)^{-4Kk}\leq\tfrac{1}{2}\lambda_{2k}$
It remains to show that $\sigma_{K}\geq\tfrac{1}{16}\sigma K$ if $\phi\leq\tfrac{1}{40}$. If $K=1$ then the claim is trivial. Suppose that $K\geq 2$. Denote $\pi_{k}=\sigma_{k}/(k\sigma)$ and $\gamma_{k}=(1+\omega^{k})/2$. Eq. (11b) can be written as $2k\sigma\pi_{2k}=2\gamma_{k}\cdot k\sigma\pi_{k}-2\phi\sigma k\sqrt{2(1-\gamma_{k})}$, therefore recursions (11) become
$\pi_{1}$ $=$ 11 11 (16a)
$\pi_{2k}$ $=$ $\gamma_{k}\pi_{k}-\phi\sqrt{2(1-\gamma_{k})}\qquad\quad\forall k\in\dot{\cal K}$ (16b)
Denoting $\Gamma_{k}=\prod_{i\in{\cal K}\cap[k]}\gamma_{i}$ and expanding these recursions gives
$\pi_{2k}=\Gamma_{k}-\phi\sum_{i\in{\cal K}\cap[k]}\tfrac{\Gamma_{k}}{\Gamma_{i}}\sqrt{2(1-\gamma_{i})}\qquad\quad\forall k\in\dot{\cal K}$ (17)
Recall that $e^{-x}>1-x>e^{-x/(1-x)}$ for $x<1$. From this we get
$\omega^{k}=(1-1/K)^{k}>e^{-k/(K-1)}>1-\tfrac{k}{K-1}\qquad\Rightarrow\qquad\gamma_{k}>1-\tfrac{k}{2(K-1)}$
We claim that $\Gamma_{K/2}\geq\tfrac{1}{8}$ [^4] Numerical evaluation suggests a better bound: $\Gamma_{K/2}>0.432$. Indeed, if $K<32$ then this can be checked numerically. Suppose that $K\geq 32$. We have $\ln(1-x)>-\tfrac{1}{1-x}\cdot x$ for $x\in(0,1)$ and thus
$\ln\Gamma_{K/2}>\sum_{k\in\dot{\cal K}}\ln\left(1-\tfrac{k}{K-1}\right)>-\tfrac{1}{1-\tfrac{16}{31}}\sum_{k\in\dot{\cal K}}\tfrac{k}{K-1}=-\tfrac{31}{15}\qquad\Rightarrow\qquad\Gamma_{K/2}\geq e^{-31/15}>\tfrac{1}{8}$
Clearly, we have $\Gamma_{k}\geq\Gamma_{K/2}>\tfrac{1}{8}$ for all $k\in\dot{\cal K}$. Therefore,
$\sum_{i\in{\cal K}\cap[K/2]}\tfrac{\sqrt{2(1-\gamma_{i})}}{\Gamma_{i}}$ $\leq 8\sum_{i\in{\cal K}\cap[K/2]}\sqrt{\tfrac{i}{K-1}}=\tfrac{8}{\sqrt{K-1}}\cdot(\sqrt{2^{0}}+\sqrt{2^{1}}+\sqrt{2^{2}}+\ldots+\sqrt{K/2})$
$=\tfrac{8}{\sqrt{K-1}}\cdot\tfrac{\sqrt{K}-1}{\sqrt{2}-1}<\tfrac{8}{\sqrt{2}-1}<20$
Using (17), we conclude that if $\phi<\tfrac{1}{40}$ then $\pi_{K}\geq\Gamma_{K/2}(1-20\phi)\geq\tfrac{1}{8}\cdot\tfrac{1}{2}=\tfrac{1}{16}$. ∎
4.1 Proof of Theorem 16
To prove the theorem, we use induction on $k\in{\cal K}$. For $k=1$ the claim is trivial. In the remainder of this section we assume that the claim holds for $k\in\dot{\cal K}$, and prove it for $2k$ and skew-symmetric matching cover ${\cal M}\subseteq{\tt Matching}$ with ${\tt size}({\cal M})\geq\delta_{2k}$. Clearly, the skew-symmetry of ${\cal M}$ implies that ${\cal M}^{k}$ is also skew-symmetric. We introduce the following notation; letter $x$ below always denotes a node in $S$.
  • Let $\mu_{x}(u,u^{\prime})$ be the expected out-degree of $x$ in $M\sim{\cal M}^{k}(u,u^{\prime})$.
  • Let $\nu_{x}(u)=\mathbb{E}_{u^{\prime}\sim_{\rho}u}[\mu_{x}(u,u^{\prime})]$ where $\rho:=\omega^{k-1}$.
  • Let ${\cal A}_{x}$ be a subset of $\mathbb{R}^{d}$ of Gaussian measure $\delta:=\tfrac{1}{2}(\delta_{2k}-\delta_{k})$ containing vectors $u$ with the largest value of $\nu_{x}(u)$, i.e. such that $\gamma_{x}:=\inf\{\nu_{x}(u)\>:u\in{\cal A}_{x}\}\geq\sup\{\nu_{x}(u)\>:u\in\mathbb{R}^{d}-{\cal A}_{x}\}$.
  • Let $\dot{\cal M}\subseteq{\cal M}$ be the matching cover obtained from ${\cal M}$ by removing edges $(x,y)$ from ${\cal M}(u)$ and edges $(y,x)$ from ${\cal M}(-u)$ for each $u\in{\cal A}_{x}$. Clearly, $\dot{\cal M}$ is a skew-symmetric matching cover with ${\tt size}(\dot{\cal M})\geq\delta_{2k}-2\delta=\delta_{k}$.
  • Let $\dot{\mu}_{x}(u,u^{\prime})$ be the expected out-degree of $x$ in $M\sim\dot{\cal M}^{k}(u,u^{\prime})$.
  • Let $\dot{\nu}_{x}(u)=\mathbb{E}_{u^{\prime}\sim_{\rho}u}[\dot{\mu}_{x}(u,u^{\prime})]$, and let $\lambda_{x}=||\dot{\nu}||_{1}$. Note that $\tfrac{1}{n}\sum_{x\in V}\lambda_{x}={\tt size}_{\rho}(\dot{\cal M}^{k})\geq\lambda_{k}$ where the last inequality is by the induction hypothesis.
Lemma 18.
$\gamma_{x}\geq\lambda_{x}$.
Proof.
By the definition of $\dot{\cal M}$, we have $\dot{\mu}_{x}(u,u^{\prime})=0$ for any $u\in{\cal A}_{x}$ and $u^{\prime}\in\mathbb{R}^{d}$. This implies that $\dot{\nu}_{x}(u)=0$ for any $u\in{\cal A}_{x}$. We have $\dot{\cal M}\subseteq{\cal M}$ and thus ˙$\dot{\cal M}^{k}\subseteq{\cal M}^{k}$. This implies that $\dot{\nu}_{x}(u)\leq\nu_{x}(u)\leq\gamma_{x}$ for any $u\in\mathbb{R}^{d}-{\cal A}_{x}$. We can now conclude that $\lambda_{x}=||\dot{\nu}_{x}||_{1}\leq\gamma_{x}\cdot 1$.
Let ${\cal H}$ be an extended matching cover where ${\cal H}(u^{\prime}_{1},u^{\prime}_{2})$ is defined as follows:
  • ()($\ast$)( ∗ )
    sample $(u^{\prime}_{1},u_{1},u_{2},u^{\prime}_{2})\sim{\cal N}_{\rho,\omega,\rho}|(u^{\prime}_{1},u^{\prime}_{2})$, sample $M_{1}\sim{\cal M}^{k}(u^{\prime}_{1},u_{1})$, sample $M_{2}\sim{\cal M}^{k}(u_{2},u^{\prime}_{2})$, compute $M=M_{1}\circ M_{2}$, output $M^{\tt good}={\tt Truncate}_{\sigma_{2k}}(M;u^{\prime}_{1})\cap{\tt Truncate}_{\sigma_{2k}}(M;u^{\prime}_{2})$.
It can be seen that ${\cal H}\subseteq{\cal M}^{2k}$ under a natural coupling. Consider the following process: sample $(u^{\prime}_{1},u^{\prime}_{2})\sim{\cal N}_{\rho\omega\rho}$ and then run procedure ()($\ast$)( ∗ ). By definitions, we have ${\tt size}_{\rho\omega\rho}({\cal H})=\tfrac{1}{n}\mathbb{E}[|M^{\tt good}|]$. Let $M^{\tt good}_{x}$ be the set of paths in $M$ that go through node $x$, and denote $\tau_{x}=\mathbb{E}[|M^{\tt good}_{x}|]$. From these definitions we get that ${\tt size}_{\rho\omega\rho}({\cal H})=\tfrac{1}{n}\sum_{x\in V}\tau_{x}$.
Lemma 19.
$\tau_{x}\geq\delta^{2/(1-\omega)}\lambda_{x}^{2}-2\varepsilon$ where $\varepsilon:=\exp\left(-\frac{\beta_{k}^{2}}{2(k+1)(1-\omega\rho)\Delta}\right)$.
Proof.
Define random variable $p_{1}$ and $p_{2}$ as follows:
  • if $x$ has an incoming path in $M_{1}$ then let $p_{1}$ be this path, otherwise let $p_{1}=\perp$;
  • if $x$ has an outgoing path in $M_{1}$ then let $p_{2}$ be this path, otherwise let $p_{2}=\perp$.
Let ${\cal E}^{\tt good}(p_{1},p_{2})=[p_{1}\neq\perp\wedge\;p_{2}\neq\perp]$. Let ${\cal E}^{\tt bad}_{1}(p_{1},u^{\prime}_{2})$ be the event that $p_{1}=(y_{1},\ldots,x)\neq\perp$, $p_{1}$ is nonviolating and $\langle x-y_{1},u^{\prime}_{2}\rangle<\sigma:=\omega\rho\sigma_{k}-\beta_{k}$. Similarly, let ${\cal E}^{\tt bad}_{2}(p_{2},u^{\prime}_{1})$ be the event that $p_{2}=(x,\ldots,y_{2})\neq\perp$, $p_{2}$ is nonviolating and $\langle y_{2}-x,u^{\prime}_{1}\rangle<\sigma$. Note, if $[{\cal E}^{\tt good}(p_{1},p_{2})\wedge\neg{\cal E}^{\tt bad}_{1}(p_{1},u^{\prime}_{2})]$ holds then $p_{1}\circ p_{2}=(y_{1},\ldots,x,\ldots,y_{2})$ is either violating or satisfies $\langle y_{2}-y_{1},u^{\prime}_{2}\rangle\geq\sigma_{k}+\sigma=\sigma_{2k}$ (and similarly for ${\cal E}^{\tt bad}_{2}(p_{2},u^{\prime}_{1})$). Therefore, $\tau_{x}\geq{\tt Pr}[{\cal E}^{\tt good}(p_{1},p_{2})]-{\tt Pr}[{\cal E}^{\tt bad}_{1}(p_{1},u^{\prime}_{2})]-{\tt Pr}[{\cal E}^{\tt bad}_{2}(p_{2},u^{\prime}_{1})]$.
Clearly, vectors $(u^{\prime}_{1},u_{1},u_{2},u^{\prime}_{2})$ are distributed according to ${\cal N}_{\rho,\omega,\rho}$. Equivalently, they are obtained by the following process: sample $(u_{1},u_{2})\sim{\cal N}_{\omega}$, sample $u^{\prime}_{1}\sim_{\rho}u_{1}$, sample $u^{\prime}_{2}\sim_{\rho}u_{2}$. Define ${\cal B}_{x}=\{u\>:\>-u\in{\cal A}_{x}\}$. By Theorem 5, we will have $(u_{1},u_{2})\in{\cal A}_{x}\times{\cal B}_{x}$ with probability at least $\delta^{2/(1-\omega)}$. Conditioned on the latter event, we have ${\tt Pr}[p_{1}\neq\perp]\geq\gamma_{x}$ and ${\tt Pr}[p_{2}\neq\perp]\geq\gamma_{x}$ (independently), where the claim for $p_{1}$ follows from the skew-symmetry of ${\cal M}^{k}$. This implies that ${\tt Pr}[{\cal E}^{\tt good}(p_{1},p_{2})]\geq\delta^{2/(1-\omega)}\gamma_{x}^{2}\geq\delta^{2/(1-\omega)}\lambda_{x}^{2}$.
We claim that ${\tt Pr}[{\cal E}^{\tt bad}_{1}(p_{1},u^{\prime}_{2})]\leq\varepsilon$. Indeed, it suffices to prove that for fixed $u^{\prime}_{1},u_{1},p_{1}$ we have ${\tt Pr}[{\cal E}^{\tt bad}_{1}(p_{1},u^{\prime}_{2})]\leq\varepsilon$ under $u_{2}\sim_{\omega}u_{1}$, $u^{\prime}_{2}\sim_{\rho}u_{2}$ (or equivalently under $u^{\prime}_{2}\sim_{\omega\rho}u_{1}$). Assume that $p_{1}=(y_{1},\ldots,x)\neq\perp$ is nonviolating (otherwise the desired probability is zero and the claim holds). Since ${\cal M}^{k}$ is $\sigma_{k}$-stretched, we have $\langle x-y_{1},u_{1}\rangle\geq\sigma_{k}$. We also have $r:=||x-y_{1}||\leq\sqrt{(k+1)\Delta}$ by Lemma 8(a). The quantity $\langle x-y_{1},u^{\prime}_{2}\rangle$ is normal with mean $\omega\rho\langle x-y_{1},u_{1}\rangle\geq\omega\rho\sigma_{k}$ and variance $(1-(\omega\rho)^{2})r^{2}$ (since e.g. we can assume that $x-y_{1}=(r,0,\ldots,0)$ by rotational symmetry, and then use the definitions of correlated Gaussians for 1-dimensional case). Therefore,
${\tt Pr}_{u^{\prime}_{2}\sim_{\omega\rho}u_{1}}\left[\langle x-y_{1},u^{\prime}_{2}\rangle<\sigma\right]$ $\leq{\tt Pr}\left[{\cal N}(\omega\rho\sigma_{k},(1-(\omega\rho)^{2})^{1/2}r)<\omega\rho\sigma_{k}-\beta_{k}\right]$
$={\tt Pr}\left[{\cal N}(0,1)<-\frac{\beta_{k}}{(1-(\omega\rho)^{2})^{1/2}r}\right]<\exp\left(-\frac{\beta_{k}^{2}}{2(1-(\omega\rho)^{2})r^{2}}\right)\leq\varepsilon$
In a similar way we prove that ${\tt Pr}[{\cal E}^{\tt bad}_{2}(p_{2},u^{\prime}_{1})]\leq\varepsilon$. The lemma follows.
We showed that
${\tt size}_{\rho\omega\rho}({\cal H})\geq\left(\tfrac{1}{n}\sum_{x\in V}\delta^{2/(1-\omega)}\lambda_{x}^{2}\right)-2\varepsilon$
Let us minimize the bound on the RHS under constraint $\tfrac{1}{n}\sum_{x\in V}\lambda_{x}\geq\lambda_{k}$. Clearly, the minimum is obtained when $\lambda_{x}=\lambda_{k}$ for all $x\in V$, in which case the bound becomes
${\tt size}_{\rho\omega\rho}({\cal H})\geq\delta^{2/(1-\omega)}\lambda_{k}^{2}-2\varepsilon\geq\lambda_{2k}$
where we used (12b) and (13).
Appendix A Matrix multiplicative weights (MW) algorithm
It this section we review the method of Arora and Kale [AK16] for the checking the feasibility of system (2) consisting of constraints (2a) and (2b)-(2e). The algorithm is given below.
1 for $t=1,2,\ldots,T$ do
2       compute
$W^{(t)}=\exp\left(\eta\sum_{r=1}^{t-1}N^{(r)}\right)\;,\quad X^{(t)}=n\cdot\frac{W^{(t)}}{{\tt Tr}(W^{(t)})}$
3       either output Fail or find “feedback matrix” $N^{(t)}$ of the form
$N^{(t)}={\tt diag}(y)+\sum_{p}f_{p}T_{p}+\sum_{S}z_{S}K_{S}-F$
where $f_{p}\geq 0$, $z_{S}\geq 0$, $\sum_{i}y_{i}+\xi n^{2}\sum_{S}z_{S}\geq\alpha$, $F$ is a symmetric matrix with $F\preceq C$, and $N^{(t)}\bullet X^{(t)}\leq 0$.
4
Algorithm 5 MW algorithm.
It can be checked that $N^{(t)}\bullet X\geq 0$ for all feasible $X$ satisfying (2) (see [AK16]). Thus, matrix $N^{(t)}$ can be viewed as a cutting plane that certifies that $X^{(t)}$ is infeasible (or lies on the boundary of the feasible region).
The procedure at line 3 is called Oracle, and the maximum possible spectral norm $||N^{(t)}||$ of the feedback matrix is called the width of the oracle. This width will be denoted as $\rho$555 Note that [AK16] formulated the algorithm in terms of the “loss matrix” $M^{(t)}=-\frac{1}{\rho}N^{(t)}$ that satisfies $||M^{(t)}||\leq 1$. Namely, it used the update $W^{(t)}\!=\!\exp\left(-\bar{\eta}\sum_{r=1}^{t-1}M^{(r)}\right)$ with $\bar{\eta}\!=\!\rho\eta$, and set $\bar{\eta}\!=\!\frac{\varepsilon}{2\rho n}$ in their Theorems 4.4 and 4.6.
Theorem 20 ([AK16, Theorems 4.4 and 4.6]).
Set $\eta=\frac{\epsilon}{2\rho^{2}n}$ and $T=\lceil\frac{4\rho^{2}n^{2}\ln n}{\epsilon^{2}}\rceil$. If Algorithm 5 does not fail during the first $T$ iterations then the optimum value of SDP (2) is at least $\alpha-\epsilon$.
The oracle used in [AK16] has the following property: if it fails then it returns a cut which is $\frac{c}{512}$-balanced and has value at most $\kappa\alpha$, where value $\kappa$ depends on the implementation. One of the implementations achieves $\kappa=O(\sqrt{\log{n}})$ and has width $\rho=\tilde{O}(\frac{\alpha}{n})$. (The runtime of this oracle will be discussed later). Setting $\epsilon=\alpha/2$ yields an $O(\kappa)=O(\sqrt{\log{n}})$ approximation algorithm that makes $\tilde{O}(1)$ calls to the oracle.
A.1 Gram decomposition and matrix exponentiation
Consider matrix $X=X^{(t)}$ computed at the $t$-th step of Algorithm 5. It can be seen that $X$ is positive semidefinite, so we can consider its Gram decomposition: $X=V^{T}V$. Let $v_{1},\ldots,v_{n}$ be the columns of $V$; clearly, they uniquely define $X$. Note that computing $v_{1},\ldots,v_{n}$ requires matrix exponentiation, which is a tricky operation because of accuracy issues. Furthermore, even storing these vectors requires $\Theta(n^{2})$ space and thus $\Omega(n^{2})$ time, which is too slow for our purposes. To address these issues, Arora and Kale compute approximations $\tilde{v}_{1},\ldots,\tilde{v}_{n}$ to these vectors using the following result.
Theorem 21 ([AK16, Lemma 7.2]).
For any constant $c>0$ there exists an algorithm that does the following: given values $\gamma\in(0,\tfrac{1}{2})$, $\lambda>0$, $\tau=O(n^{3/2})$ and matrix $A\in\mathbb{R}^{n\times n}$ of spectral norm $||A||\leq\lambda$, it computes matrix $\tilde{V}\in\mathbb{R}^{d\times n}$ with column vectors $\tilde{v}_{1},\ldots,\tilde{v}_{n}$ of dimension $d=O(\tfrac{\log n}{\gamma^{2}})$ such that matrix $\tilde{X}=\tilde{V}^{T}\tilde{V}$ has trace $n$, and with probability at least $1-n^{-c}$, one has
$|\;||\tilde{v}_{i}||^{2}-||v_{i}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}||^{2}+\tau)$ $\forall i$ (18a)
$|\;||\tilde{v}_{i}-\tilde{v}_{j}||^{2}-||v_{i}-v_{j}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}-\tilde{v}_{j}||^{2}+\tau)$ $\forall i,j$ (18b)
where $v_{1},\ldots,v_{n}$ are the columns of a Gram decomposition of $X=n\cdot\frac{\exp(A)}{{\tt Tr}(\exp(A))}$. The complexity of this algorithm equals the complexity of computing $kd$ matrix-vector products of the form $A\cdot u$, $u\in\mathbb{R}^{n}$, where $k=O(\max\{\lambda^{2},\log\frac{n^{5/2}}{\tau}\})$.
Note that matrices $A$ used in Algorithm 5 have norm at most $\eta\rho T$. Therefore, we can set $\lambda=\Theta(\frac{\rho n\log n}{\alpha})$ when applying Theorem 21 to Algorithm 5. Parameters $\gamma$ and $\tau$ will be specified later.
From now on we make the following assumption.
Assumption 1.
We have (unobserved) matrix $V\in\mathbb{R}^{n\times n}$ and (observed) matrix $\tilde{V}\in\mathbb{R}^{d\times n}$ with $X=V^{T}V$, $\tilde{X}=\tilde{V}^{T}\tilde{V}$ and ${\tt Tr}(X)={\tt Tr}(\tilde{X})=n$ satisfying conditions (18) where $v_{1},\ldots,v_{n}$ are the columns of $V$ and $\tilde{v}_{1},\ldots,\tilde{v}_{n}$ are the columns of $\tilde{V}$.
A.2 Oracle implementation
Let us denote $S=\{i\in V\>:\>||\tilde{v}_{i}||^{2}\leq 2\}$. We have $\sum_{i\in V}||\tilde{v}_{i}||^{2}={\tt Tr}(\tilde{V}^{T}\tilde{V})=n$ and thus $|S|\geq n/2$. First, one can eliminate an easy case.
Proposition 22.
Suppose that $K_{S}\bullet\tilde{X}<\tfrac{\xi n^{2}}{4}$. Then setting $y_{i}=-\tfrac{\alpha}{n}$ for all $i\in V$, $z_{S}=\tfrac{2\alpha}{\xi n^{2}}$, $z_{S^{\prime}}=0$ for all $S^{\prime}\neq S$, and $F=0$ gives a valid output of the oracle with width $\rho=O(\tfrac{\alpha}{n})$ assuming that parameters $\tau,\gamma$ in Theorem 21 satisfy $\gamma\leq\tfrac{1}{2}$ and $\tau\leq\tfrac{\xi}{2}$.
Proof.
Denote $z_{ij}=||v_{i}-v_{j}||^{2}$ and $\tilde{z}_{ij}=||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$. We know that $\tilde{Z}:=\sum_{ij}\tilde{z}_{ij}=K_{S}\bullet\tilde{X}<\tfrac{\xi n^{2}}{4}$ where the sum is over $i,j\in S$. Also, $|z_{ij}-\tilde{z}_{ij}|\leq\gamma(\tilde{z}_{ij}+\tau)$ for all $i,j$. This implies that
$K_{S}\bullet X=\sum_{ij}z_{ij}<\tilde{Z}+\gamma\tilde{Z}+\tfrac{n^{2}}{2}\gamma\tau\leq(1+\gamma)\tfrac{\xi n^{2}}{4}+\tfrac{n^{2}}{2}\gamma\tau\leq\tfrac{\xi n^{2}}{2}$
Note that $N^{(t)}=-\tfrac{\alpha}{n}I+\tfrac{2\alpha}{\xi n^{2}}K_{S}$. We have $\sum_{i}y_{i}+\xi n^{2}\sum_{S}z_{S}=n\cdot(-\tfrac{\alpha}{n})+\xi n^{2}\cdot\tfrac{2\alpha}{\xi n^{2}}=\alpha$ and $N^{(t)}\bullet X=-\tfrac{\alpha}{n}I\bullet X+\tfrac{2\alpha}{\xi n^{2}}(K_{S}\bullet X)\leq-\tfrac{\alpha}{n}\cdot n+\tfrac{2\alpha}{\xi n^{2}}\cdot\tfrac{\xi n^{2}}{2}=0$, as desired. Also, $||N^{(t)}||\leq||-\tfrac{\alpha}{n}I||+||\tfrac{2\alpha}{\xi n^{2}}K_{S}||=O(\tfrac{\alpha}{n})$. ∎
Now suppose that $\sum_{i,j\in S}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}=K_{S}\bullet\tilde{X}\geq\tfrac{\xi n^{2}}{4}$. We will set $N^{(t)}_{ij}=0$ for all $(i,j)\notin S\times S$. When describing how to set the remaining entries of $N^{(t)}$, it will be convenient to treat $N^{(t)}$, $X$, $\tilde{X}$ etc. as matrices of size $|S|\times|S|$ rather than $n\times n$, and $y$ as a vector of size $|S|$. (Note that $X$ and $\tilde{X}$ are the submatrices of the original matrices, and ${\tt Tr}(X)\leq n$). With some abuse of terminology we will redefine $V=S$ and $n=|S|$, and refer to the original variables as $V_{\tt orig}$ and $n_{\tt orig}\in[n,2n]$. Thus, from now on we make the following assumption.
Assumption 2.
$||\tilde{v}_{i}||^{2}\leq 2$ for all $i\in V$ and $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}\geq\tfrac{\xi n_{\tt orig}^{2}}{4}\geq\tfrac{\xi n^{2}}{4}$.
Let us set $y_{i}=\frac{\alpha}{n}$ for all $i\in V$ and $z_{S}=0$ for all $S$, then $\sum_{i}y_{i}+\xi n_{\tt orig}^{2}\sum_{S}z_{S}=\alpha$. Note that $N^{(t)}=\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$ and $\tfrac{\alpha}{n}I\bullet X\leq\alpha$, so condition $(\sum_{p}f_{p}T_{p}-F)\bullet X\leq-\alpha$ would imply that $N^{(t)}\bullet X\leq 0$. Our goal thus becomes as follows.
Find variables $f_{p}\geq 0$ and symmetric matrix $F\preceq C$ such that $(\sum_{p}f_{p}T_{p}-F)\bullet X\leq-\alpha$.
We arrive at the specification of Oracle given in Section 2.1.
Appendix B Proof of Theorem 5
Here we repeat the argument in [MOR+06]. For a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ denote $||f||_{p}=(\mathbb{E}_{u}[f^{p}(u)])^{1/p}$. Let $T_{\omega}f$ is the function $\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ defined via $(T_{\omega}f)(u)=\mathbb{E}_{\hat{u}\sim_{\omega}u}f(\hat{u})$. ($T_{\omega}$ is called the “Ornstein-Uhlenbeck operator”).
Theorem 23 ([Bor82]).
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$ and $-\infty<q\leq p\leq 1$, $0\leq\omega\leq\sqrt{\frac{1-p}{1-q}}$. Then
$||T_{\omega}f||_{q}\geq||f||_{p}$
Let $f,g$ be the indicator functions of sets ${\cal A},{\cal B}$ respectively. Set $p=1-\omega$ and $q=1-1/\omega$, so that $11/p+1/q=$ 1 and $\sqrt{\tfrac{1-p}{1-q}}=\omega$. Note that $||f||_{p}=||g||_{p}=\delta^{1/p}$. We can write
${\tt Pr}_{(u,u^{\prime})\sim{\cal N}_{\omega}}[(u,u^{\prime})\in{\cal A}\times{\cal B}]$ $=\mathbb{E}_{(u,u^{\prime})\sim{\cal N}_{\omega}}[f(u)g(u^{\prime})]=\mathbb{E}_{u}[fT_{\omega}g]$
$\geq||f||_{p}||T_{\omega}g||_{q}$ by reverse Hölder inequality
$\geq||f||_{p}||g||_{p}$ by Theorem 23
$=\delta^{2/(1-\omega)}$